### v1

# Information Must Flow: Recursive Bootstrapping for Information Bottleneck in Optimal Transport

Abstract
We introduce the Context-Content Uncertainty Principle (CCUP) as a unifying framework for understanding how information flows through cognitive systems by resolving the broken symmetry between high-entropy context and low-entropy content. Inference is framed as a dynamic cycle that aligns context with content through selective, bidirectional interactions, enabling information to propagate efficiently across representational hierarchies. Within this framework, cycle formation resolves the Information Bottleneck in Optimal Transport (iBOT) by coordinating bottom-up contextual disambiguation with top-down content reconstruction. This dual process is implemented via a Rao-Blackwellized inference strategy, formalized as variational entropy minimization, which steers the system toward minimal joint uncertainty while preserving the directionality of inference.
Crucially, local cycle completion gives rise to temporally extended chain formation (temporal bootstrapping) and spatially distributed hierarchical composition (spatial bootstrapping), enabling the system to simulate goals, evaluate counterfactuals, and iteratively refine internal models. To formalize this bootstrapped memory process, we prove a Delta Convergence Theorem showing that recursive entropy minimization over time yields structured memory states that converge to delta-like distributions over latent manifolds. These converged attractors act as fixed points in the flow of information, stabilizing perceptual schemas and motor plans as compact, generalizable priors. Spatial bootstrapping through ventral–dorsal functional segregation and temporal bootstrapping via the perception–action cycle support progressive refinement of sensorimotor models through continuous interaction with the environment. The sleep–wake cycle extends this dynamic across timescales, transforming episodic experiences into semantic knowledge via recursive consolidation, allowing the flow of information to bridge immediate perception with long-term abstraction.
We further propose a representational extension of CCUP in which each level of the hierarchy performs delta-seeded inference: internally generated low-entropy content seeds diffuse outward through goal-constrained inference paths shaped by top-down priors and external context. This structure-adaptive flow of information circumvents the curse of dimensionality by confining inference to task-relevant manifolds. Building on this foundation, we hypothesize that language emerges as a symbolic transport system, an operational extension of CCUP, that externalizes latent content, enabling the transfer of structured inferences across individuals and aligning intersubjective inference cycles. Together, these results establish iBOT as a foundational principle for the directed flow of information in both individual cognition and collective intelligence, situating inference itself as a structured conduit through which minds adapt, align, and extend.
Keywords Information Bottleneck in Optimal Transport (iBOT)  $\cdot$ Content-Context Uncertainty Principle (CCUP)  $\cdot$ dynamic alignment  $\cdot$ variational formation  $\cdot$ Delta convergence theorem  $\cdot$ structure-before-specificity  $\cdot$ inverted inference  $\cdot$ recursive bootstrapping
1 Problem Formulation: What is iBOT and How to Tackle it?
The Information Bottleneck (IB) is not merely a tool for supervised learning or signal compression; it is a pervasive constraint shaping biological cognition, AI, and the evolution of representational systems. At its core, the IB formalism captures a fundamental tradeoff: how to preserve task-relevant information while discarding irrelevant detail Tishby2000 . This tension manifests universally wherever a system must distill high-dimensional observations into compact internal states capable of supporting prediction, control, or generalization. In neural systems, the bottleneck appears in the selective encoding of sensory inputs Chechik2005 , the consolidation of memory Hasson2015 , and the abstraction of structure across cortical hierarchies Yamins2016 . In artificial agents, it underlies deep feature compression, latent variable modeling, and the emergence of disentangled representations Alemi2016 ; Higgins2017 . Even in natural language, perception, and planning, agents continually navigate the bottleneck between ambiguous context and structured intent Lake2017 ; Friston2010 . Recognizing the IB as a universal organizing principle allows us to reinterpret cognition itself as a process of adaptive compression, one that aligns internal structure with external demand across time, scale, and domain.
While the IB formalism captures the tension between compression and relevance, it leaves open the question of how relevant content should be transmitted or instantiated within a high-entropic context. Optimal Transport (OT) villani2009optimal provides a natural geometric solution to this question, modeling inference as the movement of probability mass from structured content $\Phi$ to observed or imagined context $\Psi$ under a cost-minimizing transport plan. However, when $\Psi$ is noisy and entangled, direct transport from context to content becomes ill-posed, mirroring the core bottleneck of inference tishby2011information . This defines what we term the Information Bottleneck in Optimal Transport (iBOT): the problem of aligning structure and specificity when transport is constrained by the need to process information. Resolving the IB in OT requires inverting the classical inference pipeline, starting from compact latent structure $\Phi$, projecting outward to generate plausible contexts $\Psi$, and refining this simulation through cycle-consistent inference. This inverted perspective transforms the bottleneck from a limitation into a guiding principle: structure-first inference allows the system to sidestep entropic overload, stabilize transport, and converge toward generalizable representations across time and scale.
The iBOT directly motivates the Context-Content Uncertainty Principle (CCUP) li2025CCUP , which frames cognition as a dynamic alignment between structured content $\Phi$ and ambiguous context $\Psi$. Whereas classical inference attempts to extract structure by compressing $\Psi$, this strategy becomes increasingly unstable in high-dimensional or underconstrained regimes. CCUP offers a structural inversion: it begins with compact, low-entropy content $\Phi$ and projects outward to simulate or interpret possible contexts $\Psi$, thereby transforming the bottleneck into an opportunity for generalization. In this view, the entropy gradient is reversed, structure precedes specificity, and inference proceeds by minimizing conditional uncertainty $H(\Psi|\Phi)$ rather than $H(\Phi|\Psi)$. This shift from bottom-up encoding to top-down simulation defines the principle of inverted inference: a generative process where structure constrains interpretation, enabling stable and scalable transport. As such, inverted inference provides a powerful mechanism for resolving iBOT by avoiding the compression of entangled context and instead organizing cognition around reusable, generalizable seeds of content.
Inverted inference gains its full expressive power when extended through recursive bootstrapping across both temporal and spatial hierarchies li2025Inverted . In the temporal domain, each content state is updated not in isolation but as part of a dynamic sequence, where structure learned at time $t$ serves as the scaffold for inference at $t+1$. This recursive alignment forms the basis of bootstrapped learning dynamics li2025CCUP , where the system iteratively refines its latent structure by minimizing conditional entropy over successive episodes. In parallel, spatial bootstrapping enables the composition of higher-level abstractions from structured subcomponents. Here, the structure is built recursively by assembling local regularities into increasingly abstract representations, satisfying both inference sufficiency and compositional coherence. The unification of these two bootstrapping processes, temporal and spatial, forms a multiscale inference engine wherein each layer constrains the next, and structure is not only inferred but grown, recursively and contextually, in space and time.
Recursive bootstrapping induces a natural convergence dynamic between content and context representations. At each iteration, the content variable $\Phi^{(t)}$ is refined to better align with the distribution of observed or simulated contexts $\Psi^{(t)}$, while simultaneously guiding the generation or interpretation of future contexts. As this process unfolds, the system iteratively reduces the conditional entropy $H(\Psi^{(t)}|\Phi^{(t)})$, concentrating probability mass around increasingly predictive latent representations. Under mild assumptions—such as continuity, compactness, and contractivity of the update operator—this dynamic converges to a fixed point $\Phi^{\ast}$, which minimizes the long-term entropy of context and satisfies the self-consistency condition $\Phi^{\ast}=\mathcal{F}(\Phi^{\ast},\Psi^{\ast})$. In the limit, this yields a delta-like representation richards1995theory : a sharply peaked distribution over latent content that collapses ambiguity while retaining a generalizable structure.
Importantly, in multi-agent settings, this recursive inference mechanism becomes socially bootstrapped: agents do not converge in isolation but through iterative alignment of their internal models via shared communicative signals. Language emerges in this framework as a population-level solution to the distributed delta-convergence problem tomasello2010origins . Rather than learning to compress content solely for internal use, agents evolve a shared latent code $Z$ that externalizes delta-converged representations, enabling others to reconstruct internal meaning with minimal uncertainty. Words, symbols, and syntactic structures become compact carriers of structured content, stabilized not through rote memorization zhang2019identity , but through repeated, entropy-aligned interactions that reinforce mutual inference. From this perspective, language is not merely a medium of expression but the emergent interface through which minds co-adapt their inference cycles, achieving compositional generalization and cycle-consistent understanding across individuals. Delta-like attractors thus arise not only within a single brain but across brains, as a product of collective abstraction grounded in socially recursive inference clark1998being .
2 Theoretical Foundation: Context–Content Uncertainty Principle (CCUP)
Cognition, at its core, is a process of resolving uncertainty robertson1929uncertainty . Every act of perception, action, or memory retrieval involves mapping ambiguous sensory or internal signals onto behaviorally relevant interpretations. We propose CCUP as a foundational framework for understanding this process. CCUP asserts that context and content are informationally asymmetric anderson1972more : although context variables may occupy a lower-dimensional sensory space, they exhibit high entropy due to ambiguity and observational limits; in contrast, content variables often span a broader semantic space but maintain low entropy due to goal-specific selection friston2005theory . CCUP characterizes inference as a directional transport process: from entropic context to structured content. This asymmetry drives the directional flow of inference - one can use context to predict content reliably, but not content to infer context unambiguously.
2.1 iBOT Calls for Dynamic Alignment of Content and Context Representation
Broken Symmetry is Ubiquitous in Information-rich Systems.
Let $\Psi$ represent high-entropy contextual variables (e.g., sensory observations or task goals) and $\Phi$ represent low-entropy content variables (e.g., object identity or action choice). The task of inference is to estimate $\Phi$ given $\Psi$, but due to the ambiguity and variability of $\Psi$, a direct mapping often leads to loss of specificity or poor generalization. For any cognitive task, an agent must infer stable latent causes from aliased sensory observations clark2013whatever . This inferential directionality, from uncertainty to specificity, is not arbitrary but structured by cognitive architecture and task demands. More fundamentally, it reflects an inherent broken symmetry in the flow of information cover1999elements , as characterized by the following lemma.
Lemma 1 (Asymmetry of Inference in Information-Rich Systems).
li2025Arrow Let $\Psi$ denote a high-entropy contextual variable, and let $\Phi$ denote a low-entropy, stable content variable inferred from $\Psi$. The CCUP formalizes an inherent asymmetry in inference between these variables, reflected in their conditional entropies. In the forward direction, the mapping $\Psi\rightarrow\Phi$ involves disambiguating aliased contextual input to infer a precise latent cause. This process is underdetermined, yielding nonzero conditional entropy: $H(\Phi|\Psi)>0$. In the reverse direction, the mapping $\Phi\rightarrow\Psi$ requires generating/simulating a range of compatible contextual instances from a compact content representation. This generalization introduces additional uncertainty: $H(\Psi|\Phi)\gg H(\Phi|\Psi)>0$.
This directional loss is not due to suboptimal learning but is structurally unavoidable given the entropy gradient between context and content. Asymmetry in inference is not merely local (directional), but is bounded globally by the mismatch in marginal entropies. Since both directions $\Psi\rightarrow\Phi$ and $\Phi\rightarrow\Psi$ carry irreducible uncertainty, the sum of these uncertainties must respect the minimum divergence in information content. Inference systems should minimize this joint uncertainty by leveraging the broken symmetry (structure prior) anderson1972more : $H(\Phi)\ll H(\Psi)$, implying that $\Phi$ should be used to condition and constrain the interpretation of $\Psi$ rao1992information ; blackwell1947conditional . We start our solution to the iBOT problem with the following lemma of entropy decomposition with a variational formulation kingma2013auto .
Lemma 2 (Entropy Decomposition Lemma with Variational Formulation).
[…]
How IB in OT Arises from the Broken Symmetry.
Under the framework of OT villani2009optimal , a high-entropic contextual distribution $\mu(\Psi)$ must be mapped onto a lower-entropic content distribution $\nu(\Phi)$, often via a latent variable $Z$. Based on CCUP, the asymmetry between $H(\Psi)$ and $H(\Phi)$ induces directional information loss. Therefore, asymmetry creates an arrow: the inference process must compensate for this directional loss by cycling- i.e., refining $\Phi$ through repeated bottom-up passes and updating $\Psi$ through top-down reconstruction. Only through such cycle formation can the system mitigate information loss and stabilize representations. Any latent encoding $\Psi\to Z$ that attempts to encode context for relevance to $\Phi$ must confront an IB: $Z$ cannot simultaneously preserve enough information for both content specificity and context recovery. Thus, the classical IB in OT is not merely a constraint on mutual information cover1999elements but a structural consequence of broken symmetry in uncertainty, which is summarized by the following lemma.
Lemma 3 (Entropy Asymmetry Implies Bottleneck in OT).
li2025Arrow Let $\Psi$ and $\Phi$ be random variables representing high-entropy context and low-entropy content, respectively. Suppose the joint distribution $p(\Psi,\Phi)$ is transported via a coupling $\gamma\in\Pi(p(\Psi),p(\Phi))$. Under CCUP, we observe: $H(\Psi)\gg H(\Phi)\Rightarrow I(\Psi;\Phi)<H(\Psi)$, which implies that any transport from $p(\Psi)$ to $p(\Phi)$ incurs an information bottleneck. Moreover, static transport mappings $T:\Psi\mapsto\Phi$ cannot preserve mutual information. To recover representational fidelity, the system must employ dynamic alignment: a cyclic process alternating between bottom-up inference $p(\Phi|\Psi)$ and top-down reconstruction $p(\Psi|\Phi)$.
The inherent asymmetry in entropy between context and content variables makes information loss during OT villani2009optimal unavoidable, creating an IB tishby2000information . Theorem 1 implies that one-way transport cannot preserve mutual information: encoding context into content inherently discards information unless explicitly corrected via reconstructive feedback. The only way to overcome this bottleneck is through dynamic alignment: an inference cycle that co-evolves content and context representations via top-down and bottom-up passes. We first give equivalent conditions for dynamic alignment below.
Definition 1 (Equivalent Conditions for Context-Content Alignment).
li2025Inverted Let $\Psi$ denote a high-entropy context variable and $\Phi$ a low-entropy content variable, with bidirectional inference maps $q(\Phi|\Psi)$ (inference) and $p(\Psi|\Phi)$ (generation). The pair $(\Psi,\Phi)$ is said to be aligned if the following equivalent conditions hold: 1) Cycle Consistency: $D_{\mathrm{KL}}(\Psi\,\|\,\hat{\Psi})\leq\varepsilon,D_{\mathrm{KL}}(\Phi\,\|\,\hat{\Phi})\leq\varepsilon$, where $\hat{\Phi}=q(\Phi|\Psi)$ and $\hat{\Psi}=p(\Psi|\hat{\Phi})$, and $\varepsilon\geq 0$ is a small alignment tolerance. 2) Mutual Information Saturation: $I(\Psi;\Phi)\approx H(\Phi)\approx H(\Psi)$, indicating that context and content share nearly all their informational degrees of freedom. 3) Entropy Symmetry (CCUP Condition): $H(\Phi|\Psi)\approx 0,H(\Psi|\Phi)\approx 0$, such that each variable fully determines the other under minimal uncertainty. 4) Optimal Transport Coupling: $\gamma^{*}=\arg\min_{\gamma\in\Pi(P_{\Psi},P_{\Phi})}\mathbb{E}_{\gamma}[c(\Psi,\Phi)]-\varepsilon H(\gamma)$, where $\gamma^{*}$ is a near-deterministic transport plan with low entropy, implying functional invertibility between context and content.
The insight underlying dynamic alignment of context and content representations informs both theoretical architectures (e.g., variational autoencoders kingma2013auto , Helmholtz machines dayan1995helmholtz ) and biological systems (e.g., ventral-dorsal loops ungerleider2000mechanisms , sleep-based memory consolidation stickgold2005sleep ), where recurrent inference cycles are necessary to counteract irreversible information loss. However, any cycle completion involves a chicken-and-egg problem, arising from the circular dependency between structured content ($\Phi$) and contextual specificity ($\Psi$).
2.2 Dynamic Alignment Requires Inverted Inference and Bootstrapping
Structure-before-Specificity Implies Inverted Inference.
In classical formulations, OT villani2009optimal faces the challenge of mapping a high-entropy contextual distribution $\mu(\Psi)$ to a lower-entropy content distribution $\nu(\Phi)$ via a latent variable $Z$. Traditional wisdom frames this as a one-shot encoding problem: how to encode $\Psi$ into $Z$ while preserving relevance to $\Phi$, which introduces an IB due to the entropy gap $H(\Psi)\gg H(\Phi)$. CCUP offers a fundamentally different perspective on introducing a feedback loop to form an asymmetric cycle based on the following observation. The broken symmetry in entropy decomposition introduces a natural directionality in inference: rather than treating $\Phi$ and $\Psi$ symmetrically, optimal inference must proceed from low-entropy structure to high-entropy specificity. In other words, the system must first identify structured content $\Phi$, and then use it to resolve ambiguity in the contextual input $\Psi$. This directional constraint gives rise to the Structure-before-Specificity (SbS) principle li2025CCUP , which we now formalize below.
Principle 1 (Structure-before-Specificity Principle).
li2025CCUP Let $\Psi$ denote high-entropy context variables and $\Phi$ denote low-entropy content variables inferred under CCUP. Let $\mathcal{D}=\{\Psi_{i}\}_{i=1}^{N}$ be a set of observed contexts drawn from an environment. Then an optimal inference strategy proceeds in two stages: 1) Structure Formation (Generalization): The agent seeks a latent structure $\Phi$ that minimizes the expected conditional entropy across contexts: $\Phi^{\ast}=\arg\min_{\Phi}\mathbb{E}_{i}\left[H(\Psi_{i}|\Phi)\right]$, which yields a generalizable representation $\Phi^{\ast}$ that captures invariant structure across variable examples; 2) Experience Binding (Specification): Once $\Phi^{\ast}$ is established, the agent performs context-specific inference by minimizing: $H(\Phi^{\ast}|\Psi_{j})$ for each new example $\Psi_{j}\in\mathcal{D}_{\text{new}}$, effectively binding content to context using the learned structural prior.
The SbS principle posits that cognitive systems achieve generalization not by overfitting to raw observations, but by grounding interpretation in structured priors that constrain what can be inferred. This principle provides a directional guide for resolving the iBOT: instead of extracting structure from noisy input, the system begins with compact representations $\Phi$ and unfolds them into context $\Psi$, selecting among possible instantiations in a manner consistent with both prior structure and observed ambiguity. To operationalize this insight, we turn to inverted inference, a computational framework in which inference is seeded not by data, but by abstract structure, and proceeds through the generation and evaluation of context-conditioned hypotheses. Inverted inference translates the SbS principle into an iterative, cycle-consistent process that uses generalization followed by specification to resolve uncertainty, dynamically aligning internal structure with external variability through recursive entropy minimization.
[…]
Figure 1: Half-Cycle Right Trick: Inverted inference starts from content $\Phi$, simulates forward to generate context $\Psi$, then re-encodes $\Psi$ to refine structure via cycle-consistency. Note that a shortcut is generated by predictive coding, which implements the updating of content without completing the cycle.
[…]
More rigorously, variational formulation introduces a recognition model $q(Z|\Psi)$ and yields an evidence lower bound (ELBO) for learning: $\log p(\Psi|\Phi)\geq\mathbb{E}_{q(Z|\Psi)}[\log p(\Psi|Z)]-D_{\mathrm{KL}}(q(Z|\Psi)\,\|\,p(Z|\Phi))$. The KL divergence term encourages cycle consistency by aligning bottom-up recognition with top-down simulation, while the reconstruction term enforces contextual predictiveness. As a result, $Z$ functions as a variational bottleneck: it absorbs ambiguity from $\Phi$ while preserving the structure necessary to generalize over possible $\Psi$. At convergence, the system achieves mutual alignment between $p(Z|\Phi)$ and $q(Z|\Psi)$, allowing recursive structure-to-context-to-structure cycles to close with minimal uncertainty. The above inverted formulation also aligns with the optimality principle of dynamic programming, which asserts that optimal substructure enables backward decomposition of global objectives bellman1966dynamic . When paired with a generative world model, goal-directed simulation becomes a powerful means of solving the search problem. It transforms the problem from open-ended expansion to a constrained inference process, where trajectories are sampled not from the full state space but from low-dimensional manifolds anchored on internally specified goals. In doing so, it turns a computational liability into an advantage: high-dimensional latent spaces offer the expressive capacity to encode rich goal hierarchies ha2018world and enable compositional planning hafner2019dream .
2.3 Bootstrapping as a Meta-Strategy for Learning and Inference
The temporal difference update rule inspires the formulation of inverted inference in Q-learning watkins1992q . Both the half-step update in Q-learning and the generalization–specification loop in inverted inference rely on the strategy of bootstrapping: using current approximations to improve future estimates in the absence of full information, thereby enabling convergence toward optimal representations through recursive refinement. The brilliant idea behind celebrated Q-learning watkins1992q is to perform a one-step bootstrapped approximation, using a partial estimate of future value to update current action values rather than computing the full trajectory to a reward. This “half-step down” idea bypasses the need for full expansion of the future, suggesting that forward search can be circumvented by local, structured inference. Q-learning sidesteps full forward search by grounding inference locally in transitions, using action-structured latent values instead of marginalizing over all futures. The trick is to bootstrap from the next state’s best action, avoiding the curse of dimensionality in planning trees bellman1966dynamic .
Just as Q-learning performs a half-step look ahead in time by using the next state’s value to update the current one, inverted inference can be interpreted as executing a half-cycle right, a generative sweep from content $\Phi$ through latent $Z$ to context $\Psi$, bootstrapping structural refinement by simulating plausible instantiations without requiring full cycle closure. Specifically, we shift the cycle formation from the original (standard inference) $\Psi\rightarrow Z\rightarrow\Phi\rightarrow Z^{\prime}\rightarrow\Psi^{\prime}$ to the right by half-cycle: $\Phi\rightarrow Z\rightarrow\Psi\rightarrow Z^{\prime}\rightarrow\Phi^{\prime}$ (inverted inference). The trick is that we do not need to complete the full cycle (the right half $\Psi\rightarrow\Phi^{\prime}$) but use the forward sweep $\Phi\rightarrow\Psi$ to refine the content representation by predictive coding keller2018predictive , which saves computation by local updates only (i.e., to update $\Phi$ from the error in $\Psi$). Formally, we have
Lemma 4 (Predictive Coding as a Half-Cycle Right Refinement).
Let $\Phi\in\mathcal{S}$ denote structured content, and let $\Psi\in\mathcal{C}$ be contextual observations. Consider the latent variable $Z\in\mathcal{Z}$ mediating the generative path from structure to context: $p(Z|\Phi),\quad p(\Psi|Z)$. Define the forward generative prediction: $\hat{\Psi}=\mathbb{E}_{Z\sim p(Z|\Phi)}\left[\mathbb{E}_{\Psi\sim p(\Psi|Z)}[\Psi]\right]$, and define the backward recognition refinement via amortized inference: $\hat{Z}=q(Z|\Psi),\quad\Phi^{\prime}=q(\Phi|\hat{Z})$. Assume the update $\Phi\to\Phi^{\prime}$ minimizes a context-aligned prediction error: $\Phi^{\prime}=\arg\min_{\Phi}D(\Psi\,\|\,\hat{\Psi}(\Phi))$, where $D$ is a divergence metric (e.g., KL or squared error). Then the half-cycle operator: $\Phi\xrightarrow{p(Z|\Phi)\rightarrow p(\Psi|Z)}\hat{\Psi}\xrightarrow{q(Z|\hat{\Psi})\rightarrow q(\Phi^{\prime}|Z)}\Phi^{\prime}$ is equivalent to a predictive coding update that refines $\Phi$ by descending the gradient of prediction error on $\Psi$.
The preceding lemma formalizes predictive coding as a half-cycle right refinement process, in which structured content $\Phi$ is updated through a generative–recognition loop grounded in prediction error minimization. This mechanism implements a local update rule that incrementally aligns simulated context $\hat{\Psi}$ with observations $\Psi$, effectively transforming forward simulation into backward structural inference. Crucially, this half-cycle framework avoids the computational burden of globally inverting high-dimensional posteriors by leveraging structured priors and amortized recognition lecun2022path . We now extend this insight to a global setting, showing that repeated application of half-cycle inference can break the curse of dimensionality by projecting high-dimensional contexts $\Psi\in\mathbb{R}^{D}$ onto low-dimensional structured representations $\Phi\in\mathbb{R}^{d}$, with $d\ll D$. Under appropriate assumptions on entropy descent and architectural regularity, this cycle-consistent bootstrapping procedure converges to a predictive fixed point $\Phi^{\ast}$ that preserves contextual information with dramatically reduced inference complexity. The following theorem formalizes this reduction in computational cost and inference dimensionality.
Theorem 1 (Breaking the Curse of Dimensionality via Half-Cycle Right Inference).
Let $\Psi\in\mathbb{R}^{D}$ denote a high-dimensional context variable, and let $\Phi\in\mathbb{R}^{d}$ denote a structured latent representation with $d\ll D$. Define a latent bottleneck variable $Z\in\mathbb{R}^{k}$ with $d\leq k\ll D$, and consider a cycle-consistent inference scheme: $\Phi^{(t)}\xrightarrow{p(Z|\Phi)}Z^{(t)}\xrightarrow{p(\Psi|Z)}\hat{\Psi}^{(t)}\xrightarrow{q(Z|\hat{\Psi})}\hat{Z}^{(t)}\xrightarrow{q(\Phi|\hat{Z})}\Phi^{(t+1)}$. Assume 1) The joint model $p(\Psi,Z,\Phi)$ factorizes through low-dimensional conditional distributions; 2) The update $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$ is entropy-decreasing: $H(\Psi|\Phi^{(t+1)})<H(\Psi|\Phi^{(t)})$; 3) The reconstruction steps $q(Z|\hat{\Psi})$, $q(\Phi|\hat{Z})$ are amortized and Lipschitz-continuous, then the inference process converges to a structured representation $\Phi^{\ast}$ that achieves predictive sufficiency: $H(\Psi|\Phi^{\ast})\approx\min_{\Phi}H(\Psi|\Phi)$ by avoiding direct marginalization over $\Psi\in\mathbb{R}^{D}$ and reducing inference complexity from $\mathcal{O}(\exp(D))$ to $\mathcal{O}(k^{2})$ under structured priors and amortization.
As an application of Theorem 1, we use two classical examples in spatial and visual inference to illustrate how half-cycle inference seeded from goal structure can break the curse of dimensionality by replacing forward trajectory expansion with predictive refinement guided by internal priors.
Example 1 (Goal-Seeded Model-based Planning in a Spatial Grid).
Consider an agent navigating a $32\times 32$ grid world, where the raw context $\Psi\in\mathbb{R}^{1024}$ encodes the full environment state (e.g., obstacle map, sensory input). The agent’s goal is to infer a path to a destination using an internal latent representation $\Phi\in\mathbb{R}^{d}$ with $d\ll 1024$, e.g., a low-dimensional code encoding goal location and path direction. Let $Z\in\mathbb{R}^{k}$ (with $d\leq k\ll D$) denote a structured latent trajectory embedding. The inference process proceeds as follows: 1) From the goal-encoded latent $\Phi^{(t)}$, simulate possible action trajectories $Z^{(t)}\sim p(Z|\Phi^{(t)})$; 2) Decode $Z^{(t)}$ into predicted environmental states $\hat{\Psi}^{(t)}\sim p(\Psi|Z^{(t)})$, e.g., simulate map states; 3) Use amortized recognition $\hat{Z}^{(t)}\sim q(Z|\hat{\Psi}^{(t)})$ and $\Phi^{(t+1)}\sim q(\Phi|\hat{Z}^{(t)})$ to update the structured goal representation. Under the assumptions of Theorem 1, this loop converges to a fixed-point representation $\Phi^{*}$ that generates predictions $\hat{\Psi}$ sufficient to capture the relevant structure of the task (i.e., $H(\Psi|\Phi^{*})\approx\min_{\Phi}H(\Psi|\Phi)$). The process avoids direct search over the $2^{1024}$-sized state space of the environment and instead bootstraps through a low-dimensional, structured latent loop with amortized inference cost $\mathcal{O}(k^{2})$.
Example 2 (Goal-Conditioned Visual Inference from Partial Observation).
Suppose a visual agent is tasked with reconstructing a full-resolution image $\Psi\in\mathbb{R}^{D}$, where $D=256\times 256\times 3=196{,}608$ represents RGB pixel intensities. Due to occlusion or partial attention, the agent only has access to partial observations of $\Psi$. The goal is to infer a plausible completion conditioned on a structured hypothesis about the scene’s content—e.g., "a dog on a couch in a living room"—encoded in a low-dimensional structured latent $\Phi\in\mathbb{R}^{d}$ with $d\ll D$. Let $Z\in\mathbb{R}^{k}$ be a generative latent code mediating between abstract structure and pixel-level predictions. In the half-cycle right inference scheme: 1) The model samples $Z^{(t)}\sim p(Z|\Phi^{(t)})$, encoding hypothesized visual attributes, textures, and layout; 2) It generates a full image prediction $\hat{\Psi}^{(t)}\sim p(\Psi|Z^{(t)})$, e.g., using a decoder trained to render photorealistic samples; 3) Using amortized inference $q(Z|\hat{\Psi}^{(t)})$ and $q(\Phi|Z)$, the system updates the structured latent $\Phi^{(t+1)}$ to reduce visual prediction error against known partial observations. Since $p(\Psi)$ lives in a prohibitively high-dimensional pixel space, directly computing $p(\Phi|\Psi)$ would require marginalizing over an intractable number of hypotheses. Theorem 1 ensures that, under entropy-decreasing updates and Lipschitz amortized inference, the model converges to a structured representation $\Phi^{*}$ such that $\hat{\Psi}=p(\Psi|Z\sim p(Z|\Phi^{*}))$ captures the missing visual information efficiently. This breaks the curse of dimensionality by projecting $\Psi$ onto a compact, compositional manifold $\Phi$, allowing generalizable reasoning over visual scenes through structure-first simulation and refinement.
3 Delta Convergence and the Emergence of Generalizable Structure
Theorem 1 demonstrates how cycle-consistent, half-cycle right inference can dramatically reduce the complexity of high-dimensional inference problems by projecting context onto structured, low-dimensional representations through amortized and entropy-aligned updates. This result highlights a broader pattern: the underlying mechanism, iterative refinement via structure-conditioned prediction and recognition hinton1995wake , is not specific to any one domain but reflects a general computational strategy. In particular, the update rule $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$ exemplifies a bootstrapping architecture in which the system uses its current estimate to simulate outcomes, then refines itself based on the discrepancy keller2018predictive . Viewed abstractly, this recursive process defines a contractive operator over a compact space of representations, with a fixed point corresponding to an optimal, self-consistent latent structure granas2003fixed . The following principle formalizes this fundamental insight: bootstrapping provides a unifying template for learning and inference as convergence to a self-consistent prediction under a surrogate refinement loop.
Principle 2 (Bootstrapping as a Unifying Principle of Learning and Inference).
Let $V_{t}\in\mathcal{X}$ be an evolving representation at iteration $t$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a compact, convex space. Suppose the update rule takes the bootstrapped form: $V_{t+1}=\mathcal{F}(V_{t},\widetilde{E}(V_{t}))$, where $\mathcal{F}:\mathcal{X}\times\mathcal{Y}\to\mathcal{X}$ is a deterministic update operator, $\widetilde{E}:\mathcal{X}\to\mathcal{Y}$ is a surrogate or predictive operator and $\mathcal{F}$ is contractive in its first argument: there exists $10<\gamma<$ 1 such that $\|\mathcal{F}(x_{1},y)-\mathcal{F}(x_{2},y)\|\leq\gamma\|x_{1}-x_{2}\|\quad\text{for all }x_{1},x_{2}\in\mathcal{X},y\in\mathcal{Y}$. Then the sequence $\{V_{t}\}_{t=0}^{\infty}$ converges to a unique fixed point $V^{\ast}\in\mathcal{X}$, satisfying: $V^{\ast}=\mathcal{F}(V^{\ast},\widetilde{E}(V^{\ast}))$.
Remark: An example of $\widetilde{E}(V_{t})$ is a partial or simulated expectation derived from $V_{t}$ itself (e.g., a predicted reward in Q-learning or a sampled context in inverted inference), and $\mathcal{F}$ is a refinement operator (e.g., Bellman update or entropy minimization). The bootstrapping principle unifies reinforcement learning sutton1998reinforcement and structure-guided inference Lake2017 as recursive fixed-point processes, where present estimates guide future refinements through partially informed simulation or projection. Bootstrapping, including half-step down and half-cycle right tricks, is computationally efficient because it turns global inference into local simulation rao1999predictive and error propagation rumelhart1985learning , reducing cost from exponential to first-order structure-preserving updates (i.e., no need for global marginalization like in the search problem minsky1961steps ).
3.1 Motivation: Fixed-point Dynamics in Bootstrapping
Bootstrapping, when framed as a meta-strategy for learning and inference, naturally gives rise to a fixed-point dynamic granas2003fixed : internal representations are recursively updated through structured feedback until a self-consistent solution emerges. This fixed-point process can be formalized as a mapping $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\Psi^{(t)})$, where each update refines structured content $\Phi$ based on the current context $\Psi$ and prior estimates. The long-term behavior of this dynamical system exhibits convergence toward delta-like attractors richards1995theory , distributions sharply peaked around highly structured, low-entropy solutions. These delta attractors represent maximally specific hypotheses that remain stable under further bootstrapped updates, serving as information-theoretic fixed points of the inference process. We start with:
Principle 3 (Iterative Convergence Principle).
Cognitive systems achieve generalization and adaptive behavior through the dynamic interplay of two complementary mechanisms: 1) Bootstrapped Learning Dynamics: Learning unfolds as a recursive, self-refining process in which current structural estimates guide the simulation and assimilation of future experiences. This enables inference and learning to proceed even under partial observability and uncertainty. 2) Delta-Like Representations: Through repeated bootstrapping, the system converges toward compact, low-entropy latent structures that serve as generalization-optimal attractors. These delta-like representations minimize both structural complexity and contextual uncertainty. Together, these mechanisms allow the brain to construct predictive, compositional models of the world that are both dynamically stable and informationally efficient.
From the perspective of CCUP, iterative convergence corresponds to the progressive alignment of ambiguous context with structured content via entropy minimization grandvalet2004semi . Such attractors are not only stable but also efficient: they encode composable abstractions that support generalization across contexts. By collapsing uncertainty into a low-dimensional manifold of self-consistent predictions, delta convergence enables both inference stability and representational reusability rigotti2013importance , key ingredients for memory formation, goal propagation, and planning under uncertainty. Formally, we have
Theorem 2 (Delta Convergence under Bootstrapped Inference).
Let $\{\Phi^{(t)}\}_{t=0}^{\infty}$ be a sequence of structured content representations evolving under a bootstrapping-based update rule: $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\Psi^{(t)})$, where $\Psi^{(t)}$ denotes the context at time $t$, and $\mathcal{F}$ is an entropy-minimizing operator that satisfies the monotonic uncertainty reduction condition: $H(\Psi^{(t)}|\Phi^{(t+1)})\leq H(\Psi^{(t)}|\Phi^{(t)})$. Assume further that $\mathcal{F}$ is contractive in a suitable metric space $(\mathcal{S},d)$ and that the context distribution stabilizes or varies within bounded entropy. Then the sequence $\{\Phi^{(t)}\}$ converges to a fixed point $\Phi^{*}$ such that: $\Phi^{*}=\lim_{t\to\infty}\Phi^{(t)}=\delta_{\phi^{*}}$, where $\delta_{\phi^{*}}$ is a delta-like distribution sharply peaked around a maximally specific latent structure $\phi^{*}\in\mathcal{S}$. Moreover, $\Phi^{*}$ minimizes the total context–content uncertainty: $\Phi^{*}=\arg\min_{\Phi}\;\mathbb{E}_{\Psi\sim P(\Psi)}\left[H(\Psi|\Phi)\right]$, and supports generalization by encoding a reusable, low-entropy structural prior consistent across multiple contexts.
The above result establishes that bootstrapped inference dynamics converge to a deterministic delta-like fixed point under a contractive and entropy-minimizing update rule. However, many real-world learning scenarios involve stochastic observations and epistemic uncertainty hullermeier2021aleatoric . In such settings, the evolution of structured content representations must be treated as a stochastic process, where each update $\Phi^{(t+1)}$ depends on sampled observations $\Psi^{(t)}\sim p(\Psi|\Phi^{(t)})$. This motivates a distributional generalization of delta convergence, in which we study the behavior of the evolving distribution $p(\Phi^{(t)})$ over structured latents richards1995theory . By extending the entropy alignment condition to the stochastic setting and assuming a weak form of contractivity, we can show that this distribution converges weakly to a delta measure concentrated at a structure $\Phi^{\ast}$ that minimizes the long-term expected uncertainty of context. The following corollary formalizes this result as a distributional analog of Theorem 2.
Corollary 1 (Delta-Like Convergence of Content under Entropy-Aligned Bootstrapping).
Let $\{\Phi^{(t)}\}\subset\mathcal{S}$ be a sequence of structured representations evolving under the recursive update: $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\Psi^{(t)}),\quad\Psi^{(t)}\sim p(\Psi|\Phi^{(t)})$, where $\mathcal{F}$ is an entropy-minimizing operator satisfying: $H(\Psi^{(t+1)}|\Phi^{(t+1)})\leq H(\Psi^{(t)}|\Phi^{(t)})$, and the update rule satisfies the contractivity condition given in Proposition 1. Then the sequence $\Phi^{(t)}$ converges to a fixed point $\Phi^{\ast}\in\mathcal{S}$, and the distribution $p(\Phi^{(t)})$ weakly converges to a delta measure: $p(\Phi^{(t)})\xrightarrow{t\to\infty}\delta(\Phi^{\ast})$, where $\Phi^{\ast}$ minimizes the long-term expected conditional entropy of context: $\Phi^{\ast}=\arg\min_{\Phi}\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}H(\Psi^{(t)}|\Phi)$.
3.2 Intuition: Generalization Benefits of Delta-like Functions
Delta-like representations emerge naturally as fixed points of recursive, entropy-aligned bootstrapping dirac1981principles . These fixed points serve as compact, predictive summaries that support generalization while minimizing contextual uncertainty. To formalize this intuition, we now establish a lemma showing that delta-like representations both minimize context–content uncertainty and constitute generalization-optimal codes under entropy-constrained inference (proof skipped).
Lemma 5 (Delta Representation as Generalization-Optimal Code).
Let $\Phi\in\mathcal{S}$ denote a structured latent variable and $\Psi\in\mathcal{C}$ a high-entropic context variable. Define the following objectives: 1) Structural complexity: $\mathcal{C}(\Phi):=H(\Phi)$; 2) Predictive specificity: $\mathcal{U}(\Phi):=H(\Psi|\Phi)$. Let the generalization-optimal representation $\Phi^{\ast}$ be defined as the minimizer of joint uncertainty: $\Phi^{\ast}=\arg\min_{\Phi}\left[H(\Phi)+\lambda H(\Psi|\Phi)\right],\quad\lambda>0$. Then the optimal solution $\Phi^{\ast}$ satisfies: 1) $p(\Phi)\to\delta(\Phi^{\ast})$ as $H(\Phi)\to 0$; 2) $\Phi^{\ast}$ minimizes contextual uncertainty: $H(\Psi|\Phi^{\ast})\leq H(\Psi|\Phi)$ for all $\Phi\in\mathcal{S}$; 3) $\Phi^{\ast}$ is the most compact structure capable of supporting maximal predictive specificity.
The preceding lemma characterizes delta representations as generalization-optimal codes that minimize both structural complexity and contextual uncertainty. These representations lie at the boundary of the information bottleneck trade-off, achieving maximal predictive specificity with minimal entropy. While the lemma establishes the optimality of such representations from a variational perspective struwe2000variational , it remains to show how they can be dynamically attained through a plausible inference process. The following theorem addresses this by demonstrating that a recursive, entropy-minimizing update rule converges to a delta-like fixed point $\Phi^{\ast}$ that satisfies the same optimality conditions. In this way, delta representations emerge not only as ideal endpoints in theory but also as stable attractors in practice through bootstrapped cycle-consistent inference li2025Inverted .
Theorem 3 (Delta Convergence Yields Generalization-Optimal Representation).
Let $\Phi^{(t)}\in\mathcal{S}\subset\mathbb{R}^{d}$ evolve under a recursive inference process: $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\Psi^{(t)}),\quad\Psi^{(t)}\sim p(\Psi|\Phi^{(t)})$, where $\mathcal{F}$ is a deterministic update operator that reduces conditional entropy: $H(\Psi^{(t+1)}|\Phi^{(t+1)})\leq H(\Psi^{(t)}|\Phi^{(t)})$, the operator $\mathcal{F}$ is contractive in $\Phi$, and the sequence $\{\Phi^{(t)}\}$ is contained in a compact, convex subset $\mathcal{S}$. Then the sequence $\Phi^{(t)}$ converges to a unique fixed point $\Phi^{\ast}\in\mathcal{S}$, and the induced distribution $p(\Phi^{(t)})$ weakly converges to a delta measure: $p(\Phi^{(t)})\xrightarrow{t\to\infty}\delta(\Phi^{\ast})$. Moreover, the fixed point $\Phi^{\ast}$ minimizes long-term contextual uncertainty: $\Phi^{\ast}=\arg\min_{\Phi\in\mathcal{S}}\mathbb{E}_{\Psi\sim p(\Psi|\Phi)}[H(\Psi|\Phi)]$, and thus defines a generalization-optimal latent representation that supports maximal predictive specificity with minimal structural complexity.
Theorem 3 demonstrates that recursive entropy-minimizing inference dynamics naturally converge to delta-like structured representations. It remains to show why such representations are desirable for learning and generalization. The following theorem establishes that delta-based codes are not merely fixed points; they are provably optimal for minimizing contextual uncertainty, generalization error, and IB complexity. This result characterizes the endpoint of bootstrapped inference as a structure that simultaneously maximizes predictive specificity and compressive efficiency.
Theorem 4 (Generalization via Delta-Based Representations).
Let $\Phi\in\mathcal{C}$ denote high-entropy contextual inputs drawn from a distribution $p(\Phi)$, and let $\Psi\in\mathcal{S}\subset\mathcal{C}$ denote low-entropy structured content representations used to guide inference over latent variables $Z$. Suppose that inference proceeds by minimizing conditional entropy $H(\Phi|\Psi)$ as prescribed by CCUP. Let $\Psi_{\delta}$ denote a delta-based representation: $p(\Psi_{\delta}|\Phi)=\delta(\Psi_{\delta}-f(\Phi))$, where $f\colon\mathcal{C}\rightarrow\mathcal{S}$ is a deterministic encoding map. Then the delta-based representation $\Psi_{\delta}$ satisfies the following: (i) It achieves minimal conditional entropy: $H(\Phi|\Psi_{\delta})\leq H(\Phi|\Psi)$ for any stochastic or smeared representation $\Psi$ with the same support. (ii) It minimizes the generalization gap under a reconstruction objective: $\mathbb{E}_{\Phi_{\text{new}}}\left[-\log p(\Phi_{\text{new}}|\Psi_{\delta})\right]\leq\mathbb{E}_{\Phi_{\text{new}}}\left[-\log p(\Phi_{\text{new}}|\Psi)\right]$, assuming $\Psi$ was trained over the same support and $\Phi_{\text{new}}\sim p(\Phi)$. (iii) It supports maximum information bottleneck compression: $I(\Phi;\Psi_{\delta})=H(\Psi_{\delta})$, since $H(\Psi_{\delta}|\Phi)=0$.
Remark: Together, these results show that delta-based representations form an optimal structural scaffold for generalization under entropy minimization and CCUP. In summary, bootstrapped convergence is not only stable but also useful for generalization li2024representations .
3.3 Variational Formulation as an Operational Principle of Inverted Inference
Having established delta convergence as the dynamical mechanism through which generalizable structure emerges, we now turn to the variational formulation of CCUP, which provides not only the underlying mathematical framework that unifies bootstrapping and delta-based inference but also an operational principle of inverted inference li2025CCUP . Specifically, CCUP defines a functional objective in which structured content $\Phi$ is optimized to minimize the expected conditional entropy of context: $\mathcal{L}(\Phi)=\mathbb{E}_{\Psi}[H(\Psi|\Phi)]$. Bootstrapping, viewed as a meta-strategy, implements a recursive update rule $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\Psi^{(t)})$ that iteratively descends this entropy landscape. Each iteration samples or simulates a context from the current content, refines the structure to reduce uncertainty, and thereby approximates variational inference under a structure-seeded prior. We consider the following model under a variational formulation.
Model 1 (Spatiotemporal Variational Inference under CCUP).
Let $\Phi_{\ell}^{(t)}$ denote the structured content at layer $\ell$, time $t$, and let $\Psi_{\ell}^{(t)}$ be the corresponding context. We define two latent variables: 1) $Z_{t}$: temporal latent variable simulating transitions from $\Phi_{\ell}^{(t)}\to\Psi_{\ell}^{(t)}$; 2) $Z_{\ell}$: spatial latent variable encoding the composition of parts $\{\Phi_{\ell-1}^{(t,i)}\}\to\Phi_{\ell}^{(t)}$. The generative model factorizes as: $p(\Psi,\Phi,Z_{t},Z_{\ell})=\prod_{t=1}^{T}\prod_{\ell=1}^{L}p(Z_{t}|\Phi_{\ell}^{(t)})\cdot p(\Psi_{\ell}^{(t)}|Z_{t})\cdot p(Z_{\ell}|\{\Phi_{\ell-1}^{(t,i)}\})\cdot p(\Phi_{\ell}^{(t)}|Z_{\ell})$. The variational posterior approximates the inverse path: $q(Z_{t},Z_{\ell}|\Psi_{\ell}^{(t)})=q(Z_{t}|\Psi_{\ell}^{(t)})\cdot q(Z_{\ell}|\Phi_{\ell}^{(t)})$. Each inference cycle minimizes the total spatiotemporal uncertainty: $\mathcal{L}=\sum_{t,\ell}\Big{[}H(\Psi_{\ell}^{(t)}|\Phi_{\ell}^{(t)})+\lambda_{\text{temp}}D_{\mathrm{KL}}(q(Z_{t}|\Psi_{\ell}^{(t)})\,\|\,p(Z_{t}|\Phi_{\ell}^{(t)}))+\lambda_{\text{space}}D_{\mathrm{KL}}(q(Z_{\ell}|\Phi_{\ell}^{(t)})\,\|\,p(Z_{\ell}|\{\Phi_{\ell-1}^{(t,i)}\}))\Big{]}$.
Remark: Latent variables $Z_{t}$ and $Z_{\ell}$ enable recursive bootstrapping in time and space, respectively. Under CCUP, this formulation minimizes contextual uncertainty while preserving structured content, guiding the convergence toward delta-like attractors across spatiotemporal inference cycles li2025Inverted . Under mild contractivity and alignment conditions, this iterative process converges to a delta-like attractor $\Phi^{*}=\delta_{\phi^{*}}$ that minimizes the CCUP objective. In this sense, delta convergence emerges as the fixed point of variational bootstrapping, encoding a reusable, low-entropy representation that supports generalization across contexts.
Lemma 6 (Delta Convergence of Structured Content).
Let $\Phi^{(t)}$ evolve under the inverted inference update: $\Phi^{(t+1)}=\arg\min_{\Phi}\left[H(\Psi|\Phi)+\lambda\,D_{\mathrm{OT}}\left(p(\Psi|\Phi),\,p(\Psi|\Phi^{(t)})\right)\right]$, Assume that: 1) The update operator is contractive in KL divergence; 2) $\Phi$ lies in a compact latent space; 3) The transport cost $D_{\mathrm{OT}}$ is continuous and convex. Then as $t\to\infty$, the distribution $p(\Phi^{(t)})\to\delta(\Phi^{*})$, and the system converges to a cycle-consistent attractor minimizing both conditional entropy and transport cost: $\Phi^{\ast}=\arg\min_{\Phi}\left[H(\Psi|\Phi)+\lambda\,D_{\mathrm{OT}}\left(p(\Psi|\Phi),\,p(\Psi|\Phi^{\ast})\right)\right]$
The preceding lemma establishes that cycle-consistent, inverted inference dynamics guided by entropy minimization and optimal transport naturally converge to a delta-like attractor. To ensure that this convergence holds more generally, across compact, convex latent spaces and under broader regularity assumptions, we formalize these dynamics in the context of the iBOT framework. The following theorem generalizes the result by showing that, under convexity and continuity of both the entropy and transport objectives, the sequence of structured representations converges to a unique fixed point $\Phi^{*}$, with the associated distribution collapsing to a delta measure centered at the generalization-optimal attractor. This result provides an operational implementation of achieving Delta convergence for iBOT.
Theorem 5 (Delta Convergence in Inverted Inference under iBOT).
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a compact, convex latent space of structured content $\Phi$. Assume: 1) The conditional entropy $H(\Psi|\Phi)$ is continuous and convex in $\Phi\in\mathcal{S}$; 2) The optimal transport cost $D_{\mathrm{OT}}(p(\Psi|\Phi),\nu)$ is continuous and convex in $\Phi$, where $p(\Psi|\Phi)$ is Lipschitz continuous in $\Phi$, and $\nu$ is a fixed target distribution over context; 3) The content update follows the inverted inference rule: $\Phi^{(t+1)}=\arg\min_{\Phi}\left[H(\Psi|\Phi)+\lambda\,D_{\mathrm{OT}}\left(p(\Psi|\Phi),\,p(\Psi|\Phi^{(t)})\right)\right]$, then the sequence $\{\Phi^{(t)}\}_{t=0}^{\infty}$ converges to a unique fixed point $\Phi^{*}\in\mathcal{S}$. Furthermore, the induced sequence of probability distributions $p(\Phi^{(t)})$ converges weakly to $\delta(\Phi^{*})$.
Remark: Theorem 5 formalizes the principle that inverted inference under the iBOT objective yields convergence toward a delta-like attractor in latent space, representing an internally stable and context-sensitive representation of structured content. This convergence underlies the brain’s capacity to compress noisy sensory streams into goal-relevant, low-entropy codes that guide perception, planning, and motor coordination. When the conditions for delta convergence are disrupted, either by noise amplification or breakdown in inference cycles, the system may exhibit pathological dynamics characterized by instability and persistent prediction errors.
To illustrate the explanatory power of this framework in a clinical setting, we now consider two examples from Parkinson’s Disease (PD) kalia2015parkinson , a disorder marked by motor tremors and impaired action control. First, we show how the tremor in PD can be understood as a failure of delta convergence, where the inference process fails to stabilize on a low-entropy motor plan. Then, we interpret focused ultrasound stimulation, a novel therapeutic approach exemplified by the work of Ali Rezai’s team bronstein2011deep , as a mechanism that restores delta convergence by externally guiding the system toward a stable attractor state, thereby suppressing tremor and reinstating precise motor control.
Example 3 (Tremor in Parkinson’s Disease as Failure of Delta Convergence).
Let $\Phi^{(t)}\in\mathcal{S}\subset\mathbb{R}^{d}$ denote a structured motor plan at iteration $t$, and let $\Psi^{(t)}\in\mathcal{C}$ represent proprioceptive or efferent feedback. In healthy motor control, inference proceeds via recursive updates: $\Phi^{(t+1)}=\arg\min_{\Phi}\left[H(\Psi^{(t)}\mid\Phi)+\lambda D_{\mathrm{KL}}(p(Z\mid\Phi)\,\|\,r(Z))\right]$, where $Z\in\mathbb{R}^{k}$ is an intermediate latent representation of motor commands, and $r(Z)$ is a prior. Under regularity conditions (e.g., contractivity of the update operator and compactness of $\mathcal{S}$), this sequence converges to a fixed point $\Phi^{\ast}$, with a sharply peaked posterior: $\lim_{t\to\infty}p(Z\mid\Phi^{(t)})=\delta(Z-Z^{\ast})$. In PD, dopaminergic degeneration impairs the precision and update dynamics of the inference loop. Let the recursive motor update be modeled as: $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$, where $\hat{\Psi}^{(t)}=\mathbb{E}_{Z\sim p(Z\mid\Phi^{(t)})}\left[\mathbb{E}_{\Psi\sim p(\Psi\mid Z)}[\Psi]\right]$ is the predicted context. Suppose that under dopamine depletion, the update operator $\mathcal{F}$ becomes non-contractive: $\|\mathcal{F}(\Phi_{1},\hat{\Psi})-\mathcal{F}(\Phi_{2},\hat{\Psi})\|\geq\|\Phi_{1}-\Phi_{2}\|\quad\text{for some }\Phi_{1},\Phi_{2}\in\mathcal{S}$; and the sequence $\{\Phi^{(t)}\}$ fails to converge. Instead, the system enters an oscillatory regime where the prediction–correction loop induces rhythmic fluctuations: $\Phi^{(t+2)}\approx\Phi^{(t)}+\epsilon\sin(\omega t),\quad\epsilon>0$, manifesting as a resting tremor. This dynamic reflects a pathological inference cycle in which prediction errors are overcorrected due to loss of damping, yielding a limit-cycle–like attractor wilson1972excitatory instead of a delta-converged motor plan. In short, tremor in PD can be interpreted as a breakdown of delta convergence in motor inference. Restorative therapies such as dopaminergic agents or deep brain stimulation may act by re-establishing contractivity in $\mathcal{F}$, allowing the system to reconverge to a stable motor goal $\Phi^{\ast}$.
Example 4 (Focused Ultrasound as Restoration of Delta Convergence in Parkinson’s Disease).
[…]
4 Spatiotemporal Bootstrapping for Fast and Slow Thinking
Inspired by the framework of slow and fast thinking kahneman2011thinking , we consider two types of bootstrapping strategies: spatial bootstrapping for fast thinking and temporal bootstrapping for slow thinking in this section. In temporal bootstrapping, the latent variable encodes a compact, goal-conditioned simulation of plausible future contexts given current structure, enabling forward rollouts in episodic memory tulving2002episodic . In spatial bootstrapping, the latent variable encodes a compositional binding pattern over lower-level parts, mediating the construction of higher-order structure treisman1980feature . In both cases, $Z$ acts as an intermediate scaffold that facilitates entropy-aligned inference while preserving structural coherence.
Definition 2 (Spatiotemporal Bootstrapping for iBOT).
Let $\Phi_{\ell}^{(t)}\in\mathcal{S}_{\ell}$ denote the structured content representation at layer $\ell$ and time $t$, and $\Psi_{\ell}^{(t)}\in\mathcal{C}_{\ell}$ its associated contextual input. The brain forms predictive models through recursive bootstrapping in both time and space:
1) Temporal Bootstrapping (Episodic Memory): At each time step $t$, content $\Phi^{(t+1)}$ is refined via cycle-consistent updates seeded by past structure and generalized through simulated context: $\Phi^{(t+1)}=\mathcal{F}_{\mathrm{time}}(\Phi^{(t)},\Psi^{(t)}),\quad\Psi^{(t)}\sim p(\Psi|\Phi^{(t)})$, with the update minimizing conditional entropy: $H(\Psi^{(t)}|\Phi^{(t+1)})<H(\Psi^{(t)}|\Phi^{(t)})$.
2) Spatial Bootstrapping (Hierarchical Composition): Within each layer $\ell$, higher-level content $\Phi_{\ell}$ is constructed by composing lower-level components $\{\Phi_{\ell-1}^{(i)}\}$, enabling part–whole generalization: $\Phi_{\ell}=\mathcal{F}_{\mathrm{space}}\left(\Phi_{\ell-1}^{(1)},\ldots,\Phi_{\ell-1}^{(n_{\ell})}\right)$, with the update minimizing spatial conditional entropy: $H(\Psi_{\ell}|\Phi_{\ell})<\sum_{i=1}^{n_{\ell}}H(\Psi_{\ell-1}^{(i)}|\Phi_{\ell-1}^{(i)})$.
Remark: Both temporal and spatial bootstrapping induce convergence toward delta-like attractors: $\Phi_{\ell}^{(\infty)}=\delta(\Phi_{\ell}^{\ast})$, which minimizes joint uncertainty: $\Phi^{\ast}=\arg\min_{\Phi}\left[H(\Phi)+\lambda H(\Psi|\Phi)\right]$. The above definition provides a unifying framework for understanding dual-process cognition as a tradeoff between spatial and temporal bootstrapping of inference. In this view, System 1 corresponds to fast thinking, which arises from spatial bootstrapping: previously learned structured content $\Phi$ is rapidly reused to resolve ambiguous context $\Psi$ through pattern completion ramsauer2020hopfield . This inference is efficient, feedforward, and compositional, minimizing conditional entropy $H(\Psi|\Phi)$ without extensive refinement. In contrast, System 2 reflects slow thinking, driven by temporal bootstrapping, in which $\Phi$ is iteratively refined through recursive cycles of structure-content alignment. This process is formalized as an inverted inference update that minimizes both conditional entropy and transport cost frogner2015learning : $\Phi^{(t+1)}=\arg\min_{\Phi}\left[H(\Psi|\Phi)+\lambda\,D_{\mathrm{OT}}\left(p(\Psi|\Phi),\,p(\Psi|\Phi^{(t)})\right)\right]$. Here, deliberate reasoning emerges as a temporally extended inference trajectory that incrementally resolves uncertainty in context by simulating plausible latent structures alemi2016deep . CCUP thus explains dual-process cognition as arising from a dynamic balance between the reuse of known structure and the generation of novel structure in response to contextual ambiguity genevay2018learning . In this section, we separately study temporal and spatial bootstrapping for slow and fast thinking next.
4.1 Episodic Memory Supports Goal-Directed Simulation via Temporal Bootstrapping
Episodic memory provides a temporal scaffold for constructing internally generated simulations that guide future behavior schacter2008episodic . Rather than functioning as a static repository of past events, episodic memory supports goal-directed inference by incrementally refining structural expectations over time bar2009proactive . Each remembered episode contributes a noisy, partial context $\Phi^{(t)}$, which is integrated into a cumulative predictive code $\Psi^{(t)}$. This code serves as a latent summary of temporally extended experience and is recursively updated to minimize uncertainty about future contextual inputs schapiro2017complementary . Crucially, this bootstrapping process does not require full marginalization over the high-dimensional posterior space but instead propagates structure through local updates (thanks to the half-cycle right trick) that balance predictive specificity and temporal continuity. As the inference unfolds, the representation $\Psi^{(t)}$ converges toward a compact, generalizable attractor $\Psi^{\ast}$, enabling the system to simulate plausible outcomes under uncertainty wang2018prefrontal . The following theorem formalizes this process as a convergence result for temporally bootstrapped inference.
Theorem 6 (Delta Convergence of Temporal Bootstrapping).
Let $\{\Psi^{(t)}\}_{t=1}^{\infty}\subset\mathcal{P}(\mathcal{S})$ be a sequence of probability distributions over structured representations, updated recursively via: $\Psi^{(t+1)}=\arg\min_{\Psi}\left[H(\Phi^{(t)}|\Psi)+\lambda D_{\mathrm{KL}}(\Psi\,\|\,\Psi^{(t)})\right]$, where $\Phi^{(t)}$ is a high-entropy contextual observation at time $t$, and $\lambda>0$ is a regularization parameter enforcing continuity. Assume: 1) (Stationarity) The contextual distribution $p(\Phi^{(t)})$ is stationary; 2) (Entropy gap) There exists a unique global minimizer $\Psi^{\ast}\in\mathcal{S}$ such that: $\Psi^{\ast}=\arg\min_{\Psi}\ \mathbb{E}_{\Phi}[H(\Phi|\Psi)]$ and for all $\Psi\neq\Psi^{\ast}$, $H(\Phi|\Psi)>H(\Phi|\Psi^{\ast})$; 3) (Sufficient smoothness) The entropy function $H(\Phi|\Psi)$ is convex and differentiable in $\Psi$, then the sequence $\Psi^{(t)}$ converges weakly to a delta distribution centered at $\Psi^{\ast}$: $\lim_{t\to\infty}\Psi^{(t)}=\delta_{\Psi^{\ast}}$ in the weak-topology on $\mathcal{P}(\mathcal{S})$. Moreover, $\lim_{t\to\infty}H(\Phi|\Psi^{(t)})=H(\Phi|\Psi^{\ast}),\quad\lim_{t\to\infty}\mathrm{Var}(\Psi^{(t)})=0$.
The convergence of episodic bootstrapping toward a generalizable latent code hinges on the underlying dynamics of cycle-consistent refinement kumaran2016learning . In particular, when each update to the structured content $\Phi^{(t)}$ is informed by a prediction $\hat{\Psi}^{(t)}$ decoded from latent variables and aligned with observed context, the system effectively implements an entropy-minimizing cycle grandvalet2004semi . These iterative updates form a contractive process in the latent space, progressively sharpening the correspondence between structure and context. When the entropy of the context given structure $H(\Psi|\Phi^{(t)})$ is strictly decreasing and bounded below, the system converges to a unique attractor $\Phi^{\ast}$ that supports stable, delta-like predictions. The following corollary formalizes this convergence behavior, demonstrating how delta-like representations naturally emerge from the interplay of entropy minimization and cycle-consistent inference.
Corollary 2 (Delta-like Convergence from Entropy-Minimizing Cycles).
Under the assumptions of Theorem 1, suppose the cycle-consistent update $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$ satisfies: 1) $\Phi^{(t)}\in\mathcal{S}\subset\mathbb{R}^{d}$ with $d\ll D$, and 2) The entropy sequence $H(\Psi|\Phi^{(t)})$ is strictly decreasing and bounded below. Then there exists a unique fixed point $\Phi^{\ast}\in\mathcal{S}$ such that: $\Phi^{\ast}=\mathcal{F}(\Phi^{\ast},\hat{\Psi}^{\ast}),\quad\text{with}\quad\hat{\Psi}^{\ast}\sim p(\Psi|Z^{\ast}),\quad Z^{\ast}\sim p(Z|\Phi^{\ast})$, and $\lim_{t\to\infty}H(\Psi|\Phi^{(t)})=H(\Psi|\Phi^{\ast})\approx 0$. Hence, $\Phi^{\ast}$ behaves as a delta-like representation: $p(\Psi|\Phi^{\ast})\approx\delta(\Psi-\hat{\Psi}^{\ast})$, defining a generalization-optimal, cycle-consistent latent code that supports maximal predictive specificity with minimal structural complexity.
4.2 Hierarchical Composition as Spatial Bootstrapping
Hierarchical composition functions as a spatial bootstrapping mechanism in cognition lee2003hierarchical . At each layer $\ell$, a structured content representation $\Phi_{\ell}$ is constructed by recursively composing and abstracting over lower-level parts $\{\Phi_{\ell-1}^{(i)}\}$. This enables complex perceptual categories, such as objects, scenes, or symbols, to be inferred from simpler components through reusable structure and alignment, rather than being learned tabula rasa dicarlo2012does . As with episodic memory in time, spatial bootstrapping allows the brain to generalize efficiently across space by iteratively building structure from previously grounded elements eichenbaum2014time . We start with the construction of a new model for spatiotemporal bootstrapping.
Model 2 (Spatiotemporal Variational Inference under CCUP).
Let $\Phi_{\ell}^{(t)}$ denote the structured content at layer $\ell$, time $t$, and let $\Psi_{\ell}^{(t)}$ be the corresponding context. We define two latent variables: 1) $Z_{t}$: temporal latent variable simulating transitions from $\Phi_{\ell}^{(t)}\to\Psi_{\ell}^{(t)}$; 2) $Z_{\ell}$: spatial latent variable encoding the composition of parts $\{\Phi_{\ell-1}^{(t,i)}\}\to\Phi_{\ell}^{(t)}$. The generative model factorizes as: $p(\Psi,\Phi,Z_{t},Z_{\ell})=\prod_{t=1}^{T}\prod_{\ell=1}^{L}p(Z_{t}|\Phi_{\ell}^{(t)})\cdot p(\Psi_{\ell}^{(t)}|Z_{t})\cdot p(Z_{\ell}|\{\Phi_{\ell-1}^{(t,i)}\})\cdot p(\Phi_{\ell}^{(t)}|Z_{\ell})$. The variational posterior approximates the inverse path: $q(Z_{t},Z_{\ell}|\Psi_{\ell}^{(t)})=q(Z_{t}|\Psi_{\ell}^{(t)})\cdot q(Z_{\ell}|\Phi_{\ell}^{(t)})$. Each inference cycle minimizes the total spatiotemporal uncertainty: $\mathcal{L}=\sum_{t,\ell}\Big{[}H(\Psi_{\ell}^{(t)}|\Phi_{\ell}^{(t)})+\lambda_{\text{temp}}D_{\mathrm{KL}}(q(Z_{t}|\Psi_{\ell}^{(t)})\,\|\,p(Z_{t}|\Phi_{\ell}^{(t)}))+\lambda_{\text{space}}D_{\mathrm{KL}}(q(Z_{\ell}|\Phi_{\ell}^{(t)})\,\|\,p(Z_{\ell}|\{\Phi_{\ell-1}^{(t,i)}\}))\Big{]}$.
The preceding model formalizes how spatiotemporal structure can be disentangled into separate but coupled latent paths: temporal inference through $Z_{t}$, simulating goal-conditioned transitions over time, and spatial inference through $Z_{\ell}$, composing lower-layer features into increasingly abstract structured codes whittington2022disentanglement . Crucially, both forms of inference obey the CCUP principle li2025CCUP : minimizing contextual uncertainty via structured priors and cycle-consistent refinement. When this bootstrapped inference is applied recursively across hierarchical levels, each layer incrementally compresses context into more structured, less ambiguous representations. The result is a progressive sharpening of the conditional distribution $p(\Phi_{\ell-1}|\Psi_{\ell})$, steering inference toward compact, generalizable representations. The following theorem formalizes this convergence in the limit of deep hierarchical composition, showing that inference over contextually grounded latent variables converges to a delta-like distribution concentrated around an abstract structural manifold.
Theorem 7 (Delta Convergence of Hierarchical Composition).
Let $\{\Psi_{\ell}\}_{\ell=0}^{L}$ be a hierarchy of latent variables forming a compositional structure under the conditions of Theorem LABEL:thm:HCSS. Suppose: 1) Each level $\Psi_{\ell}$ is obtained via entropy-aligned composition: $\Psi_{\ell}=\arg\min_{\Psi}\left[H(\Phi_{\ell-1}|\Psi)+\lambda_{\ell}D_{\mathrm{KL}}\left(\Psi\,\|\,f_{\ell}(\Psi_{\ell-1}^{(1)},\ldots,\Psi_{\ell-1}^{(n_{\ell})})\right)\right]$; 2) The conditional entropy decreases at each level: $H(\Phi_{\ell-1}|\Psi_{\ell})<H(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)}),\quad\forall i$; 3) The KL divergence across levels is uniformly bounded: $D_{\mathrm{KL}}\left(p(\Phi_{\ell-1}|\Psi_{\ell})\,\|\,p(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)})\right)\leq\epsilon_{\ell}$, then, in the limit $\ell\to L$, the top-level representation $\Psi_{L}$ converges in distribution to a delta-like measure concentrated around a minimal uncertainty manifold $\mathcal{M}^{\ast}\subset\mathcal{S}_{L}$, such that: $\lim_{\ell\to L}p(\Phi_{\ell-1}|\Psi_{\ell})\to\delta_{\mathcal{M}^{\ast}}(\Phi_{\ell-1})$ This implies that inference becomes maximally specific with respect to context, while remaining globally structured and compositionally abstract.
Theorem 7 establishes that hierarchical context representations $\Psi_{\ell}$ converge in distribution toward delta-like manifolds of minimal uncertainty, this phenomenon also implies a deep structural constraint on the content representations $\Phi_{\ell}$ that generate them. Because each context $\Psi_{\ell}$ is synthesized through a generative bottleneck conditioned on $\Phi_{\ell}$, the sharpening of contextual predictions feeds back to recursively refine the content representations themselves. This yields a bootstrapped alignment between context-specific structure and content-specific abstraction across layers and time. The following theorem formalizes this dual convergence dolgopolik2024convergence : under entropy-decreasing generative cycles and differentiable composition, content representations $\Phi_{\ell}$ progressively collapse onto delta-like attractors that act as minimal sufficient statistics for context generation. These fixed points form a hierarchy of compositional abstractions, each layer capturing structure at a broader scope and lower entropy, converging toward generalization-optimal codes.
Theorem 8 (Hierarchical Delta-Convergence of Content Representations).
Let $\Phi_{\ell}^{(t)}\in\mathbb{R}^{d_{\ell}}$ denote the structured content at layer $\ell$ and time $t$, with $d_{\ell}\ll D$, and let $\Psi_{\ell}^{(t)}\in\mathbb{R}^{D}$ denote the corresponding context at that layer. Assume: 1) Each inference cycle follows: $\Phi_{\ell}^{(t+1)}=\mathcal{F}_{\ell}\left(\Phi_{\ell}^{(t)},\hat{\Psi}_{\ell}^{(t)}\right),\quad\hat{\Psi}_{\ell}^{(t)}\sim p(\Psi_{\ell}|Z_{\ell}),\quad Z_{\ell}\sim p(Z_{\ell}|\Phi_{\ell}^{(t)})$, where $Z_{\ell}$ is a latent variable bottleneck at layer $\ell$. 2) For each $\ell$, the conditional entropy is strictly decreasing and bounded below: $H(\Psi_{\ell}^{(t)}|\Phi_{\ell}^{(t)})>H(\Psi_{\ell}^{(t+1)}|\Phi_{\ell}^{(t+1)})\geq 0$. 3) Each layer composes its representation from the layer below: $\Phi_{\ell}^{(t)}=f_{\ell}\left(\Phi_{\ell-1}^{(t,1)},\dots,\Phi_{\ell-1}^{(t,n_{\ell})}\right)$, for some differentiable composition operator $f_{\ell}$, then for each layer $\ell$, the sequence $\{\Phi_{\ell}^{(t)}\}_{t=1}^{\infty}$ converges to a delta-like fixed point $\Phi_{\ell}^{\ast}\in\mathbb{R}^{d_{\ell}}$, such that: $\Phi_{\ell}^{\ast}=\mathcal{F}_{\ell}\left(\Phi_{\ell}^{\ast},\hat{\Psi}_{\ell}^{\ast}\right),\quad p(\Psi_{\ell}|\Phi_{\ell}^{\ast})\approx\delta\left(\Psi_{\ell}-\hat{\Psi}_{\ell}^{\ast}\right)$. Hence, hierarchical inference produces a tower of delta-like attractors: $\{\Phi_{\ell}^{\ast}\}_{\ell=1}^{L}$, each represents a generalization-optimal scaffold for context prediction at increasing levels of abstraction and spatial composition.
Remark: Delta-converged latent codes unify memory consolidation and symbolic abstraction as complementary forms of representational grounding in cognitive systems. In summary, the framework of spatiotemporal bootstrapping offers a unifying account of fast and slow thinking as emergent properties of context-sensitive inference: spatial bootstrapping enables rapid generalization through structural reuse, while temporal bootstrapping supports deliberative reasoning through recursive refinement. This dual-process architecture reflects the brain’s need to balance efficiency with flexibility across timescales.
Corollary 3 (Delta-Convergence as the Basis of Cognitive Grounding).
Let $\{\Phi_{\ell}^{(t)}\}_{t=1}^{\infty}$ be the structured representations produced by recursive inference at layer $\ell$ and time $t$, governed by the entropy-minimizing update: $\Phi_{\ell}^{(t+1)}=\mathcal{F}_{\ell}\left(\Phi_{\ell}^{(t)},\hat{\Psi}_{\ell}^{(t)}\right),\quad\hat{\Psi}_{\ell}^{(t)}\sim p(\Psi_{\ell}|Z_{\ell}),\quad Z_{\ell}\sim p(Z_{\ell}|\Phi_{\ell}^{(t)})$. Suppose: 1) Temporal bootstrapping (memory): Over episodes $t$, the sequence $\Phi_{\ell}^{(t)}$ converges to a fixed point: $\lim_{t\to\infty}\Phi_{\ell}^{(t)}=\Phi_{\ell}^{\ast},\quad\text{with}\quad p(\Psi_{\ell}|\Phi_{\ell}^{\ast})\approx\delta(\Psi_{\ell}-\hat{\Psi}_{\ell}^{\ast})$. 2) Spatial bootstrapping (composition): Each $\Phi_{\ell}^{\ast}$ is composed from lower-layer attractors: $\Phi_{\ell}^{\ast}=f_{\ell}(\Phi_{\ell-1}^{(1)\ast},\dots,\Phi_{\ell-1}^{(n_{\ell})\ast})$, and satisfies a fixed-point composition constraint: $\Phi_{\ell}^{\ast}=\mathcal{F}_{\ell}(\Phi_{\ell}^{\ast},\hat{\Psi}_{\ell}^{\ast})$, Then each $\Phi_{\ell}^{\ast}$ defines a grounded cognitive representation: either as memory (a temporally stable, generalization-optimal schema) or as symbol (a spatially compositional, context-invariant scaffold).
Remark: These findings suggest that delta-converged latent codes play a critical role in grounding both memory consolidation and symbolic abstraction, enabling cognitive systems to generalize efficiently while refining representations over time. The dual mechanisms of spatial and temporal bootstrapping thus serve as complementary engines of inference: the former scaffolds immediate recognition and generalization through reusable structural motifs, while the latter enables gradual refinement of beliefs through recursive, cycle-consistent updating. Together, they instantiate the brain’s adaptive balance between fast and slow thinking, leveraging compressed structure to navigate a high-dimensional and uncertain world. When these bootstrapping mechanisms break down, either due to failure in maintaining latent structure or disruption of cyclic refinement, the result can be catastrophic for cognitive function. In the next two examples, we illustrate this vulnerability by examining frontotemporal dementia (FTD) as a clinical manifestation of broken inverted inference, where abstract goal representations can no longer organize perceptual and semantic content. Conversely, we show how targeted memory restoration interventions can be understood as reestablishing cycle-consistent inference dynamics, thereby reinstating the alignment between structured content and contextual variability necessary for coherent memory retrieval and flexible cognition.
Example 5 (Frontotemporal Dementia as a Breakdown of Inverted Inference).
Frontotemporal dementia (FTD) manifests as a progressive degeneration of the frontal and/or temporal lobes, leading to cognitive and behavioral impairments including loss of planning, disinhibition, semantic deficits, and emotional blunting bang2015frontotemporal . We interpret these symptoms as a consequence of disrupted recursive bootstrapping within the inverted inference framework. Let $\Phi\in\mathcal{S}$ represent structured internal content (e.g., semantic memory or social norms), and let $\Psi^{\text{obs}}_{t}\in\mathcal{C}$ denote contextual observations across time. In a healthy system, inference proceeds via the following generative-recursive loop: $Z_{t}\sim P(Z\mid\Phi),~{}\hat{\Psi}_{t}\sim P(\Psi\mid Z_{t}),~{}Z_{t}^{\prime}\sim Q(Z\mid\Psi^{\text{obs}}_{t}),~{}\hat{\Phi}_{t}\sim P(\Phi\mid Z_{t}^{\prime})$. This loop enables predictive coding and cycle-consistent refinement of structure: $\Phi^{(t+1)}\leftarrow\arg\min_{\Phi}\mathcal{D}\left(\Psi^{\text{obs}}_{t}\,\|\,P(\Psi\mid Z_{t})\right)$, where $\mathcal{D}$ is a divergence metric, and $Z_{t}\sim P(Z\mid\Phi^{(t)})$. In FTD, distinct breakdowns occur depending on the locus of degeneration: 1) Frontal atrophy impairs temporal bootstrapping by degrading the update dynamics $\Phi^{(t)}\mapsto\Phi^{(t+1)}$, leading to inflexible behavior and loss of goal-directed reasoning; 2) Temporal atrophy disrupts the generative prior $P(Z\mid\Phi)$, resulting in degraded simulations $\hat{\Psi}_{t}$ and collapsed semantic content; 3) Global breakdown of the inference cycle prevents alignment between predicted and observed context, leading to poor self-monitoring, semantic incoherence, and social dysfunction. In short, FTD can be viewed as a failure of inverted inference: the system can no longer simulate, match, and refine latent structure $\Phi$ through recursive prediction-error minimization. This offers a computational explanation for the disintegration of narrative self, semantic memory, and adaptive behavior in FTD.
Example 6 (Memory Restoration as Recovery of Cycle-Consistent Inference).
[…]
5 Social Bootstrapping and the Origin of Language
From Spatiotemporal Bootstrapping to Social Bootstrapping.
While spatiotemporal bootstrapping enables individual agents to construct coherent internal models by integrating information across scales of space (e.g., object composition) and time (e.g., episodic memory), such inference remains fundamentally bounded by the agent’s own sensory history and computational constraints. To transcend these individual limitations, cognitive systems engage in social bootstrapping clark1998being , a process through which structured latent content is externalized, shared, and iteratively refined via communicative interaction. In any act of communication, a speaker must transform a rich internal state, comprising beliefs, intentions, emotions, or memories, into a compressed external signal such as language, gesture, or expression tomasello2010origins . This transformation inevitably involves information loss, posing a fundamental tradeoff: how can the speaker reduce the complexity of the message while still preserving its relevance to the listener’s interpretive task? This tradeoff is captured by the IB principle, which formalizes the problem of compressing internal content in a way that maximally preserves task-relevant information. In the context of social communication, the IB framework models how agents encode and decode signals to align internal models under constraints of noise, ambiguity, and limited bandwidth. It provides a powerful lens for understanding how language, emotion, and shared context evolve to mitigate communicative uncertainty and enable intersubjective inference.
Definition 3 (Information Bottleneck in Social Communication).
Let $\Phi\in\mathcal{S}$ denote the structured internal state (e.g., latent intent, belief, or concept) of a speaker, and let $\Psi\in\mathcal{C}$ be the observable signal (e.g., utterance, gesture, expression) transmitted during communication. Let $\hat{\Phi}\in\mathcal{S}$ denote the listener’s inferred internal state based on observing $\Psi$. The Information Bottleneck for Social Communication is the problem of finding a stochastic encoding $p(\Psi|\Phi)$ that minimizes the mutual information between the internal state and the signal, $\min_{p(\Psi|\Phi)}I(\Phi;\Psi)$, subject to a constraint on the relevance of the signal to the listener’s inference task: subject to $\text{subject to }I(\Psi;\hat{\Phi})\geq\mathcal{I}_{0}$, where $\mathcal{I}_{0}$ is a lower bound on the amount of relevant information that must be preserved for successful communication. This tradeoff captures the core challenge of social communication: compressing internal content into efficient signals that remain maximally informative under the listener’s reconstruction process.
In social contexts, this mechanism becomes the basis for social bootstrapping, a process whereby intersubjective constraints recursively anchor individual inference cycles. Through imitation, joint attention, and communicative feedback, agents co-construct shared latent models, refining their internal representations via socially mediated error correction clark2015surfing . Just as episodic memory bootstraps inference through time, social bootstrapping extends the cycle across minds, creating a distributed inference network grounded in a shared structure. In this extended framework, language emerges as a socially bootstrapped code that enables the transmission of compressed internal content across agents macwhinney1998models , aligning distributed inference cycles through shared priors and mutual simulation. Thus, the principles of cycle-consistent bootstrapping scale from the internal architecture of the individual to the emergent coordination of collective intelligence, transforming private inference loops into intersubjective bridges that support knowledge accumulation, cultural transmission, and cooperative reasoning boyd2005origin .
Building on this foundation, we hypothesize that language emerges as a symbolic transport system, an operational extension of CCUP that externalizes latent content, enabling structured inferences to be transferred across individuals jackendoff2003precis . Words, phrases, and syntactic constructions act as compact carriers of internal structure, preserving delta-compressed representations and guiding listener inference along goal-aligned paths mercer2002words . In this view, language is not merely a vehicle for communication, but a cognitive infrastructure for synchronizing inference cycles, a solution to the collective binding problem posed by intersubjective cognition scholte2025beyond . Formally, we have
Lemma 7 (Language as a Latent Variable Solving IB for OT in Social Communication).
Let $\Phi\in\mathcal{S}$ be the structured internal content of a speaker (e.g., belief, intention), and $\Psi\in\mathcal{C}$ be the observable communicative signal (e.g., speech, gesture). Let $Z\in\mathcal{L}$ denote a shared latent linguistic representation. Language acts as an intermediating variable $Z$ that solves the Information Bottleneck (IB) for Optimal Transport (OT) in social communication, by satisfying the following: 1) Compression: The map $\Phi\to Z$ minimizes the encoding cost while preserving communicative relevance: $\min_{p(Z|\Phi)}I(\Phi;Z)\quad\text{subject to}\quad I(Z;\hat{\Phi})\geq\mathcal{I}_{0}$; 2) Transport: The signal $\Psi$ is generated from $Z$ in a way that minimizes semantic transport mismatch: $\min_{p(\Psi|Z)}D_{\mathrm{OT}}\big{(}p(\Psi|Z),\,p(\Psi|\hat{Z})\big{)}$; 3) Cycle-consistency: The listener infers $\hat{Z}\to\hat{\Phi}$, and successful communication occurs when $\hat{\Phi}\approx\Phi$, closing the inference loop. Hence, language defines a shared latent space $\mathcal{L}$ that compresses, transports, and reconstructs structured meaning across agents.
[…]
(a) Single-Agent IB–OT Triangle
$\Phi_{1}$$\Phi_{2}$$\Phi_{3}$hared Latent Language $Z\in\mathcal{L}$$\Psi_{1\to 2}$$\Psi_{2\to 3}$$\Psi_{3\to 1}$ulti-Agent Alignment via Language
(b) Social Bootstrapping Across Agents
Figure 2: iBOT in Social Communication: (a) single-agent inference (Language emerges as an intermediating latent variable $Z$); (b) multi-agent bootstrapping via shared latent language codes (Social bootstrapping enables emergence and refinement of the shared latent variable $Z$ through interaction among agents).
From Individual Communication to Collective Alignment.
The preceding formulation characterizes language as a structured latent variable $Z$ that enables efficient and cycle-consistent communication between a single speaker and listener by solving the tradeoff between compression and semantic transport. However, natural language does not arise in isolation; it emerges, stabilizes, and evolves within a community of interacting agents. As communication scales from dyadic exchanges to multi-agent systems, the pressure to align divergent internal states across agents drives the emergence of a shared latent codebook mercer2002words . In this broader setting, language becomes a socially bootstrapped scaffold for collective inference, iteratively refined through feedback, imitation, and mutual prediction. The following result formalizes this transition by showing how shared symbolic codes emerge as optimal solutions to a multi-agent iBOT problem, where each agent seeks to compress and broadcast its internal state in a way that minimizes inference error across the population. We have the following result regarding the emergent property of language.
Theorem 9 (Emergent Language as a Shared Latent Variable for Multi-Agent Alignment).
Let $\{\Phi_{i}\in\mathcal{S}\}_{i=1}^{N}$ denote the internal latent states of $N$ communicating agents. Let $Z\in\mathcal{L}$ be a shared latent linguistic codebook, and let $\Psi_{i\to j}\in\mathcal{C}$ denote the signal emitted by agent $i$ and received by agent $j$. The social system solves the following collective optimization problem: $\min_{\{p(Z|\Phi_{i}),\,p(\Psi_{i\to j}|Z)\}}\sum_{i\neq j}\left[I(\Phi_{i};Z)+\lambda\,D_{\mathrm{OT}}\big{(}p(\Psi_{i\to j}|Z),\,p(\Psi_{i\to j}|\hat{Z}_{j})\big{)}\right]$ subject to: $\mathbb{E}_{i\neq j}\left[I(Z;\hat{\Phi}_{j})\right]\geq\mathcal{I}_{0}$, where each agent reconstructs $\hat{Z}_{j}\to\hat{\Phi}_{j}$, and successful communication occurs when $\hat{\Phi}_{j}\approx\Phi_{i}$. Social bootstrapping enforces shared priors on latent linguistic codebook $Z$, enabling convergence to a cycle-consistent and socially grounded latent codebook.
From Shared Alignment to Symbolic Structure.
Together, th insights in Lemma 7 and Theorem 9 establish the iBOT as a foundational principle for directed information flow in both individual cognition and collective intelligence. Inference itself becomes a structured conduit through which minds adapt, align, and extend, recursively anchoring structure to context across space, time, and social interaction labov1970study . Establishing a shared latent codebook allows a population of agents to reliably exchange internal states via compressed, interpretable signals. However, mere alignment is not sufficient for scalability and generalization: as the semantic space grows in complexity, the communication system must also acquire an internal structure to support compositional reuse, abstraction, and productivity peirce1995social . In this setting, social bootstrapping not only facilitates convergence to a common code but also promotes the emergence of symbolic structure, where latent codes factorize into disentangled components, each mapping to a specific semantic dimension mordatch2018emergence . The next lemma formalizes this process, showing that under natural pressures such as communicative reward, entropy minimization, and transport efficiency, the shared codebook evolves toward a compositional architecture that reflects the structured geometry of latent meaning.
Lemma 8 (Emergence of Symbolic Compositionality via Social Bootstrapping).
Let $\Phi_{i}\in\mathcal{S}$ be the internal latent states of communicating agents, and let $Z\in\mathcal{L}$ be a shared latent linguistic code evolving under repeated social interactions. Suppose agents update their encoders $p(Z|\Phi_{i})$ and decoders $p(\hat{\Phi}_{j}|Z)$ to minimize: $\mathcal{L}=I(\Phi_{i};Z)+\lambda\,D_{\mathrm{OT}}(p(\Psi|Z),p(\Psi|\hat{Z}))-\eta\,R(\hat{\Phi}_{j}\approx\Phi_{i})$, where $R$ is a communicative reward for successful inference. Under the assumptions of shared priors across agents, smooth differentiable updates (e.g., reinforcement learning, Hebbian plasticity), and a pressure toward communication efficiency (low entropy signals), the codebook $Z\in\mathcal{L}$ evolves toward a compositional structure: $Z\cong z_{1}\otimes z_{2}\otimes\cdots\otimes z_{k}$, where each $z_{i}$ encodes a disentangled factor of variation in $\Phi$, enabling compositional generalization and symbolic reuse.
From Compositional Structure to Delta Convergence
The emergence of a compositional codebook reflects a population-level adaptation toward efficient and generalizable communication. In such systems, meanings are expressed by combining discrete symbolic units, words, gestures, or morphemes, each corresponding to an interpretable factor of latent content. This compositionality enables systematic generalization across a combinatorial semantic space kirby2015compression ; lake2018generalization . However, compositional structure alone is not sufficient for robust coordination: without stable symbol–meaning mappings, communication becomes noisy, ambiguous, or inconsistent. To support cycle-consistent inference across individuals, the mappings between latent factors and symbolic components must converge reliably across agents, time, and usage contexts, which calls for a refinement process that goes beyond structure and seeks convergence.
Specifically, we model each component of the symbolic code as undergoing an iterative update shaped by three forces: (i) communicative feedback, which rewards successful reconstruction and penalizes misalignment; (ii) entropy minimization, which compresses ambiguous symbol-to-meaning distributions into confident, peaked ones; and (iii) transport alignment, which regularizes changes to ensure semantic continuity over time. These forces act together to sharpen the conditional distribution $p(\phi_{i}|z_{i})$, for each symbolic slot $z_{i}$, such that the posterior over meanings collapses to a delta-like distribution centered on a distinct semantic value richards1995theory . This delta convergence transforms a fuzzy, redundant symbolic system into one that is discrete, interpretable, and compositionally reliable. Symbol–meaning associations stabilize across communicative exchanges, enabling agents to bootstrap consistent inference cycles. As a result, the symbolic codebook becomes not only compositional in form, but also delta-convergent in function, anchoring abstract meaning in minimal-entropy representations that support generalization, alignment, and cultural transmission mordatch2018emergence . Such attractors are not artifacts of overfitting, but the natural consequence of recursive social inference operating under constraints of entropy and transport cost. The following theorem establishes sufficient conditions under which this delta convergence emerges as the optimal solution to the compositional inference problem.
Theorem 10 (Delta Convergence of Symbolic Composition).
Let $\Phi\in\mathcal{S}\subset\mathbb{R}^{d}$ denote a structured latent meaning, and let $Z=(z_{1},\dots,z_{k})\in\mathcal{L}$ be a compositional code with $k$ discrete symbol slots (e.g., words, morphemes, gesture primitives). Define an iterative update scheme for each code component: $z_{i}^{(t+1)}=\arg\min_{z_{i}}\left[H(\Phi|z_{i})+\lambda\,D_{\mathrm{OT}}(p(\Phi|z_{i}),\,p(\Phi|z_{i}^{(t)}))\right]$ Assume: 1) Each update operator is contractive in KL divergence; 2) $\mathcal{L}_{i}$ is a finite discrete set (symbol vocabulary for slot $i$); 3) The meaning space $\mathcal{S}$ decomposes as $\Phi=(\phi_{1},\dots,\phi_{k})$, with conditional independence $p(\Phi|Z)\approx\prod_{i}p(\phi_{i}|z_{i})$; 4) Agents receive communicative feedback that enforces inference cycle-consistency: $\Phi\to Z\to\hat{\Phi}\approx\Phi$. Then as $t\to\infty$, each symbolic component $z_{i}^{(t)}$ converges to a delta-like attractor: $p(z_{i}^{(t)})\to\delta(z_{i}^{\ast})$ and the full language code converges to a compositional fixed point: $Z^{\ast}=(z_{1}^{\ast},\dots,z_{k}^{\ast}),\quad\text{minimizing}\quad\sum_{i}\left[H(\phi_{i}|z_{i})+\lambda D_{\mathrm{OT}}(\cdot)\right]$ Therefore, language emerges as a symbolic system whose structure reflects the factorization of latent meaning under joint entropy-transport minimization.
Finally, we demonstrate the application of Theorem 10 with the following two examples. The first illustrates how discrete symbolic components converge to delta-like assignments that encode distinct semantic dimensions. The second explains how all human languages, while diverse in form, reflect delta-convergent solutions to the problem of compositional communication.
Example 7 (Toy Language for Object Description).
Consider a simple communication task in which agents must describe objects composed of two latent factors: color and shape. Let the latent space be $\Phi=(\phi_{\text{color}},\phi_{\text{shape}})$, where $\phi_{\text{color}}\in\{\texttt{red},\texttt{blue},\texttt{green}\}$ and $\phi_{\text{shape}}\in\{\texttt{circle},\texttt{square},\texttt{triangle}\}$. The compositional code $Z=(z_{1},z_{2})$ consists of two discrete symbols: one assigned to color ($z_{1}$) and one to shape ($z_{2}$). Each symbol is selected from a finite vocabulary $\mathcal{L}_{1}=\mathcal{L}_{2}=\{a,b,c\}$. Agents interact repeatedly: one selects an object $\Phi=(\phi_{\text{color}},\phi_{\text{shape}})$, encodes it into a message $Z=(z_{1},z_{2})$, and the listener decodes $Z$ to produce an inferred object $\hat{\Phi}$. After each episode, agents update their symbol assignments $z_{i}^{(t)}$ to minimize: $H(\Phi|z_{i})+\lambda D_{\mathrm{OT}}(p(\Phi|z_{i}),p(\Phi|z_{i}^{(t)}))$ subject to feedback $\hat{\Phi}\approx\Phi$. Initially, symbol use is noisy and inconsistent (e.g., $a$ may refer to either red or green), resulting in ambiguous communication. Over time, delta convergence ensures that each symbol stabilizes to a specific semantic value: $p(z_{1}=a)\to\delta(\texttt{red}),\quad p(z_{2}=b)\to\delta(\texttt{circle})$ leading to a stable compositional codebook such as: $z_{1}\in\{a\mapsto\texttt{red},\;b\mapsto\texttt{blue},\;c\mapsto\texttt{green}\},\quad z_{2}\in\{a\mapsto\texttt{circle},\;b\mapsto\texttt{square},\;c\mapsto\texttt{triangle}\}$. This example illustrates how discrete symbolic components converge to delta-like assignments that encode distinct semantic dimensions, allowing agents to generalize to novel combinations (e.g., blue triangle) through compositional reuse.
Example 8 (Cross-Linguistic Convergence in Motion Event Encoding).
Let $\Phi=(\phi_{\text{manner}},\phi_{\text{path}},\phi_{\text{goal}})\in\mathcal{S}$ denote a structured latent meaning corresponding to a motion event, such as “The girl ran into the house.” This semantic tuple includes: 1) $\phi_{\text{manner}}$: the manner of motion (e.g., running); 2) $\phi_{\text{path}}$: the trajectory (e.g., entering); 3) $\phi_{\text{goal}}$: the spatial endpoint (e.g., the house). Let $Z=(z_{1},z_{2},z_{3})\in\mathcal{L}_{1}\times\mathcal{L}_{2}\times\mathcal{L}_{3}$ be a compositional symbolic representation of $\Phi$, where each $z_{i}$ is selected from a finite vocabulary slot corresponding to one semantic factor. Across languages, different mappings from $\Phi\to Z\to\Psi$ emerge, but each converges to a delta-like attractor within its symbolic system: 1) English (Satellite-framed): $z_{1}=\texttt{run}$, $z_{2}=\texttt{into}$, $z_{3}=\texttt{the house}$Rightarrow “The girl ran into the house.”
[…]
Despite these typological differences, each language exhibits: 1) Factorization: a mapping from latent semantics $\Phi$ to symbolic slots $Z$; 2) Stability: repeated use leads to convergence to discrete symbol choices $z_{i}^{\ast}$; 3) Expressive equivalence: the capacity to communicate the same structured meaning across systems. This supports the conclusion of Theorem 10: the symbolic code $Z$ in each language converges to a fixed point minimizing joint entropy and transport distortion over semantic structure.
6 Conclusion: Information Must Flow
This paper has advanced a unified framework for understanding the directional flow of information in cognitive and communicative systems, grounded in the interplay between the IB, OT, and the CCUP. We began by identifying a fundamental constraint—termed the inference bottleneck of transport (iBOT), that arises from the need to compress high-dimensional internal states into low-bandwidth signals while preserving their relevance for inference. Overcoming iBOT requires organizing inference not as a symmetric mapping between signals and meanings, but as a directional flow that aligns structured content with contextual constraints, in accordance with CCUP.
At the level of individual cognition, this challenge is addressed through inverted inference, which reverses the traditional data-driven flow by initiating inference from structured, low-entropy goals. This directional process is further refined through spatiotemporal bootstrapping, whereby content is constructed across nested hierarchies of spatial composition and temporal integration. These mechanisms collectively solve iBOT by enabling internal models to emerge through recursive alignment between structured priors and context-sensitive updates.
Extending this framework to collective intelligence introduces additional complexity: when multiple agents aim to align latent meanings via communication, a new intermediating variable is needed. We identified language as this variable, a socially bootstrapped symbolic code that enables the externalization and reconstruction of structured content across minds. Language resolves the multi-agent form of iBOT by providing a shared latent space through which context and content can be jointly optimized. Crucially, the compositional structure of language, its ability to systematically combine symbols to express complex meanings, emerges as a natural consequence of Delta convergence, wherein repeated social feedback and inference cycles drive symbolic codes toward discrete, disentangled, and reusable representations.
Together, these results demonstrate that information must flow, not arbitrarily, but directionally, guided by principles that balance compression, transport, and inference. From the recursive architecture of individual cognition to the symbolic scaffolds of social communication, overcoming iBOT is the central computational challenge that shapes the evolution of intelligent systems.
Appendix A Proof of Theorem 1
Proof.
We aim to show that the half-cycle right inference procedure described by the update
$\Phi^{(t)}\xrightarrow{p(Z|\Phi)}Z^{(t)}\xrightarrow{p(\Psi|Z)}\hat{\Psi}^{(t)}\xrightarrow{q(Z|\hat{\Psi})}\hat{Z}^{(t)}\xrightarrow{q(\Phi|\hat{Z})}\Phi^{(t+1)}$
converges to a structured representation $\Phi^{\ast}$ that minimizes the predictive uncertainty $H(\Psi|\Phi)$, while reducing the computational cost of inference.
Step 1: Factorization and Dimension Reduction. By assumption, the joint distribution factorizes as: $p(\Psi,Z,\Phi)=p(\Phi)p(Z|\Phi)p(\Psi|Z)$, with each component involving conditional distributions over variables in lower-dimensional spaces: $\Phi\in\mathbb{R}^{d}$, where $d\ll D$; and $Z\in\mathbb{R}^{k}$, where $d\leq k\ll D$. This implies that the model can generate $\Psi\in\mathbb{R}^{D}$ through a two-stage low-dimensional latent pathway: $\Phi\mapsto Z\mapsto\Psi$. Thus, inference via $\Phi\mapsto Z\mapsto\Psi$ avoids directly marginalizing over $\Psi\in\mathbb{R}^{D}$, reducing the effective complexity of inference from exponential in $D$ to polynomial in $k$.
Step 2: Monotonic Decrease in Entropy. By assumption, the inference update satisfies: $H(\Psi|\Phi^{(t+1)})<H(\Psi|\Phi^{(t)})$, which implies that the sequence $\{H(\Psi|\Phi^{(t)})\}$ is strictly decreasing and bounded below by zero. Hence, by the monotone convergence theorem, the sequence converges:
$H(\Psi|\Phi^{(t)})\to H(\Psi|\Phi^{\ast})\quad\text{as }t\to\infty.$
Therefore, $\Phi^{(t)}\to\Phi^{\ast}$ in the space of structured representations, and $\Phi^{\ast}$ minimizes conditional entropy.
Step 3: Convergence Under Lipschitz and Amortization Assumptions. Let us define the update operator $\mathcal{F}\colon\mathbb{R}^{d}\times\mathbb{R}^{D}\to\mathbb{R}^{d}$ via: $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$, where $\hat{\Psi}^{(t)}\sim p(\Psi|Z^{(t)})$, $Z^{(t)}\sim p(Z|\Phi^{(t)})$. Since both $q(Z|\hat{\Psi})$ and $q(\Phi|\hat{Z})$ are assumed to be amortized and Lipschitz-continuous, and the composite operator $\mathcal{F}$ is entropy-decreasing, Banach’s fixed-point theorem ensures that the iterative update converges to a unique fixed point $\Phi^{\ast}$ in $\mathbb{R}^{d}$.
Step 4: Computational Advantage. The naive computation of $p(\Phi|\Psi)\propto p(\Psi|\Phi)p(\Phi)$ via marginalization over $\Psi\in\mathbb{R}^{D}$ would require evaluating or integrating over an exponential number of configurations (e.g., $\mathcal{O}(\exp(D))$). Instead, the proposed inference scheme propagates through $Z\in\mathbb{R}^{k}$ with amortized recognition and decoding. Assuming structured priors (e.g., Gaussian or autoregressive forms), the forward–backward latent updates can be computed in $\mathcal{O}(k^{2})$, which is exponentially more efficient for high-dimensional $D$.
Conclusion. The cycle-consistent half-step inference path converges to a fixed point $\Phi^{\ast}$ that minimizes $H(\Psi|\Phi)$ and achieves approximate predictive sufficiency. By avoiding full inversion of $\Psi\in\mathbb{R}^{D}$ and exploiting structured low-dimensional priors and amortized updates, the method breaks the curse of dimensionality by reducing computational complexity from $\mathcal{O}(\exp(D))$ to $\mathcal{O}(k^{2})$. ∎
Appendix B Proof of Theorem 2
Proof.
We proceed in three parts.
Step 1 (Monotonic entropy descent). By assumption, the operator $\mathcal{F}$ satisfies: $H(\Psi^{(t)}|\Phi^{(t+1)})\leq H(\Psi^{(t)}|\Phi^{(t)})$, for each $t\geq 0$. This forms a non-increasing sequence of conditional entropies bounded below by zero: $H(\Psi^{(0)}|\Phi^{(0)})\geq H(\Psi^{(1)}|\Phi^{(1)})\geq\cdots\geq 0$. Hence, the sequence converges: $\lim_{t\to\infty}H(\Psi^{(t)}|\Phi^{(t)})=H^{*}\geq 0$.
Step 2 (Contractive mapping convergence). Since $\mathcal{F}$ is contractive in the complete metric space $(\mathcal{S},d)$, i.e.,
$d(\mathcal{F}(\Phi_{1},\Psi),\mathcal{F}(\Phi_{2},\Psi))\leq\lambda d(\Phi_{1},\Phi_{2}),\quad\text{for all }\Phi_{1},\Phi_{2}\in\mathcal{S},\text{ with }0<\lambda<1,$
it follows from the Banach Fixed-Point Theorem that the sequence $\{\Phi^{(t)}\}$ converges to a unique fixed point $\Phi^{*}\in\mathcal{S}$ such that: $\Phi^{*}=\mathcal{F}(\Phi^{*},\Psi^{*})$, where $\Psi^{*}:=\lim_{t\to\infty}\Psi^{(t)}$ (assumed to converge or fluctuate within bounded entropy). Hence, the fixed point $\Phi^{*}$ is a stable attractor under recursive inference updates.
Step 3 (Characterization as delta-like entropy minimizer). By convergence and monotonicity of conditional entropy, we have: $H(\Psi^{*}|\Phi^{*})=\min_{\Phi}H(\Psi^{*}|\Phi)$. In the context of inference, this corresponds to minimizing the expected uncertainty of context given structured content. The entropy is minimized when the conditional distribution $P(\Psi|\Phi)$ is sharply peaked, i.e., when $\Phi$ deterministically predicts $\Psi$. This occurs when $\Phi^{*}\approx\delta_{\phi^{*}}$, where $\phi^{*}\in\mathcal{S}$ is the maximally informative structural latent that organizes multiple observations $\Psi$.
Therefore, $\Phi^{*}$ minimizes total expected context–content uncertainty: $\Phi^{*}=\arg\min_{\Phi}\mathbb{E}_{\Psi\sim P(\Psi)}\left[H(\Psi|\Phi)\right]$, and acts as a reusable, low-entropy structural prior supporting generalization across contexts. ∎
Appendix C Proof of Theorem 3
Proof.
We aim to show that the sequence $\{\Psi^{(t)}\}\subset\mathcal{P}(\mathcal{S})$ defined by the iterative update:
$\Psi^{(t+1)}=\arg\min_{\Psi\in\mathcal{P}(\mathcal{S})}\left[H(\Phi^{(t)}|\Psi)+\lambda D_{\mathrm{KL}}(\Psi\,\|\,\Psi^{(t)})\right]$
converges weakly to a delta distribution $\delta_{\Psi^{\ast}}$, where $\Psi^{\ast}$ uniquely minimizes the conditional entropy $\mathbb{E}_{\Phi}[H(\Phi|\Psi)]$.
Step 1: Objective Function and Gradient Flow. Define the composite objective functional:
$\mathcal{L}(\Psi;\Phi^{(t)},\Psi^{(t)}):=H(\Phi^{(t)}|\Psi)+\lambda D_{\mathrm{KL}}(\Psi\,\|\,\Psi^{(t)}).$
Since $H(\Phi|\Psi)$ is convex and differentiable in $\Psi$ by assumption, and KL divergence is strictly convex in $\Psi$, the full objective is strictly convex. Therefore, $\Psi^{(t+1)}$ exists and is unique at each time step.
Step 2: Contractivity of the Update. Let $\Psi^{\ast}$ be the unique global minimizer of the stationary functional: $\Psi^{\ast}=\arg\min_{\Psi}\mathbb{E}_{\Phi}[H(\Phi|\Psi)]$, and let us define the distance function: $\mathcal{D}^{(t)}:=D_{\mathrm{KL}}(\Psi^{\ast}\,\|\,\Psi^{(t)})$. From Pinsker’s inequality and the stationarity of $p(\Phi^{(t)})$, we can bound the entropy gap: $H(\Phi^{(t)}|\Psi^{(t)})-H(\Phi|\Psi^{\ast})\geq c\cdot\|\Psi^{(t)}-\Psi^{\ast}\|_{1}^{2}$ for some $c>0$, as long as $\Psi^{(t)}\neq\Psi^{\ast}$. Furthermore, the update minimizes a Bregman-like proximal step around $\Psi^{(t)}$, which satisfies: $D_{\mathrm{KL}}(\Psi^{\ast}\,\|\,\Psi^{(t+1)})<D_{\mathrm{KL}}(\Psi^{\ast}\,\|\,\Psi^{(t)})$ whenever $\Psi^{(t)}\neq\Psi^{\ast}$, due to strict convexity and the entropy gap.
[…]
Step 4: Collapse to Delta Measure. By assumption, $\Psi^{\ast}$ is the minimizer of the entropy functional and is unique. Since the update is constrained to probability distributions, and the minimum of a convex functional over a convex domain occurs at an extremal point when the objective is strictly convex, the limit must collapse onto a delta distribution centered at $\Psi^{\ast}$: $\lim_{t\to\infty}\Psi^{(t)}=\delta_{\Psi^{\ast}}$, in the weak-* topology. As the distribution converges to a delta function, its variance vanishes: $\lim_{t\to\infty}\mathrm{Var}(\Psi^{(t)})=0$.
Step 5: Entropy Convergence. From continuity of entropy in weak-* topology under bounded support (since $\mathcal{S}$ is compact), $\lim_{t\to\infty}H(\Phi|\Psi^{(t)})=H(\Phi|\Psi^{\ast})$.
Appendix D Proof of Theorem 4
Proof.
Let $\Phi\in\mathcal{C}$ be a high-entropy contextual input variable, and let $\Psi\in\mathcal{S}\subset\mathcal{C}$ be a structured latent variable derived from $\Phi$ by an encoder $f\colon\mathcal{C}\to\mathcal{S}$, such that: $p(\Psi_{\delta}|\Phi)=\delta(\Psi_{\delta}-f(\Phi))$. This defines a deterministic representation $\Psi_{\delta}=f(\Phi)$. We now verify each of the three claims.
(i) Minimal Conditional Entropy. By definition of conditional entropy, $H(\Phi|\Psi)=\mathbb{E}_{\Psi}\left[H(p(\Phi|\Psi))\right]$. If $\Psi$ is stochastic (e.g., with $p(\Psi|\Phi)$ spread over a region), then there is a nonzero residual uncertainty in recovering $\Phi$ from $\Psi$. In contrast, for the delta encoding $\Psi_{\delta}=f(\Phi)$, the joint distribution is: $p(\Phi,\Psi_{\delta})=p(\Phi)\cdot\delta(\Psi_{\delta}-f(\Phi))$. Then, $H(\Phi|\Psi_{\delta})=\mathbb{E}_{\Psi_{\delta}}\left[H(p(\Phi|\Psi_{\delta}))\right]$. Since $\Psi_{\delta}$ is a deterministic function of $\Phi$, we can invert $f$ in the support of $\Phi$, making the conditional distribution $p(\Phi|\Psi_{\delta})$ a pushforward under $f^{-1}$. Therefore, the mapping $\Phi\mapsto\Psi_{\delta}$ does not add any uncertainty, only possibly collapses redundant information. So we have $H(\Phi|\Psi_{\delta})\leq H(\Phi|\Psi)$, with equality if and only if $\Psi$ also deterministically specifies $\Phi$, which proves (i).
(ii) Minimal Generalization Gap. Consider the expected log-loss on unseen data: $\mathbb{E}_{\Phi_{\text{new}}}\left[-\log p(\Phi_{\text{new}}|\Psi)\right]$. For delta-based encoding, we have $\Psi_{\delta}=f(\Phi_{\text{new}})$, and assuming the same decoder (e.g., Bayes-optimal or variational), the conditional distribution $p(\Phi|\Psi_{\delta})$ minimizes the cross-entropy term since $\Psi_{\delta}$ is a minimal sufficient statistic under the CCUP principle. By the data processing inequality and sufficiency of $\Psi_{\delta}$, we know:
$\mathbb{E}_{\Phi_{\text{new}}}\left[-\log p(\Phi_{\text{new}}|\Psi_{\delta})\right]\leq\mathbb{E}_{\Phi_{\text{new}}}\left[-\log p(\Phi_{\text{new}}|\Psi)\right].$
Thus, delta codes minimize the generalization error over latent representations, proving (ii).
(iii) Maximum Information Bottleneck Compression. The mutual information between $\Phi$ and $\Psi_{\delta}$ is: $I(\Phi;\Psi_{\delta})=H(\Psi_{\delta})-H(\Psi_{\delta}|\Phi)$. But since $\Psi_{\delta}=f(\Phi)$, a deterministic function, we have $H(\Psi_{\delta}|\Phi)=0,\quad\text{so}~{}I(\Phi;\Psi_{\delta})=H(\Psi_{\delta})$. Hence, delta-based representations achieve the maximum possible compression under the information bottleneck principle for a given $H(\Psi_{\delta})$, proving (iii).
Conclusion. Taken together, (i)–(iii) show that delta-based representations: - minimize contextual uncertainty, - generalize optimally across unseen contexts, - and serve as minimal sufficient codes, all consistent with CCUP’s principle of entropy-aligned structure–context inference.
Appendix E Proof of Theorem 5
Proof.
[…]
Appendix F Proof of Theorem 6
Proof.
We aim to show that the sequence $\{\Psi^{(t)}\}_{t=1}^{\infty}\subset\mathcal{P}(\mathcal{S})$, evolving via the entropy-regularized update: $\Psi^{(t+1)}=\arg\min_{\Psi}\left[H(\Phi^{(t)}|\Psi)+\lambda D_{\mathrm{KL}}(\Psi\,\|\,\Psi^{(t)})\right]$, converges weakly to a delta distribution centered at the unique generalization-optimal representation $\Psi^{\ast}$.
Step 1: Convex Objective Guarantees Uniqueness. Define the functional: $\mathcal{L}(\Psi;\Phi^{(t)},\Psi^{(t)}):=H(\Phi^{(t)}|\Psi)+\lambda D_{\mathrm{KL}}(\Psi\,\|\,\Psi^{(t)})$. By assumption (3), the first term $H(\Phi^{(t)}|\Psi)$ is convex and differentiable in $\Psi$. The KL divergence is strictly convex in $\Psi$ over the interior of the probability simplex. Therefore, $\mathcal{L}$ is strictly convex, and each update step admits a unique minimizer $\Psi^{(t+1)}$.
Step 2: Descent Property and Boundedness. Let $H^{*}:=H(\Phi|\Psi^{\ast})$. From assumption (1), the distribution of $\Phi^{(t)}$ is stationary. Taking expectations over $p(\Phi^{(t)})=p(\Phi)$, we have:
$\mathbb{E}[H(\Phi^{(t)}|\Psi^{(t+1)})]+\lambda D_{\mathrm{KL}}(\Psi^{(t+1)}\,\|\,\Psi^{(t)})\leq\mathbb{E}[H(\Phi|\Psi^{(t)})].$
This shows that the expected contextual uncertainty $H(\Phi|\Psi^{(t)})$ is non-increasing and bounded below by $H^{*}$, so: $\lim_{t\to\infty}H(\Phi|\Psi^{(t)})=H^{\infty}\geq H^{*}$.
Step 3: Convergence to Minimizer. Assumption (2) states that $\Psi^{\ast}$ is the unique minimizer of $\mathbb{E}[H(\Phi|\Psi)]$, and any deviation from $\Psi^{\ast}$ strictly increases $H(\Phi|\Psi)$. If $\Psi^{(t)}\not\to\Psi^{\ast}$, there would be infinitely many $t$ for which: $H(\Phi|\Psi^{(t)})>H^{*}+\epsilon$, for some $\epsilon>0$, contradicting monotonic convergence toward $H^{\infty}=H^{*}$. Hence, $\Psi^{(t)}\to\Psi^{\ast}$.
Step 4: Weak-* Convergence to Delta. Since $\Psi^{(t)}\to\Psi^{\ast}$, and $\Psi^{\ast}\in\mathcal{S}$ is a point-mass minimizer under the entropy objective, we conclude: $\Psi^{(t)}\xrightarrow{w^{*}}\delta_{\Psi^{\ast}}$, i.e., convergence in the weak-* topology.
Step 5: Collapse of Variance and Entropy. Because convergence is to a delta distribution, it follows that:
$\lim_{t\to\infty}\mathrm{Var}(\Psi^{(t)})=0,\quad\lim_{t\to\infty}H(\Phi|\Psi^{(t)})=H(\Phi|\Psi^{\ast}).$
Conclusion. The recursive entropy-regularized updates form a contraction toward the unique minimizer $\Psi^{\ast}$, and the KL regularization ensures stability. Therefore, $\Psi^{(t)}\to\delta_{\Psi^{\ast}}$ in the weak-* sense, and the sequence minimizes both contextual uncertainty and variance, completing the proof. ∎
Appendix G Proof of Theorem 7
Proof.
We proceed by induction over the hierarchy levels $\ell=1,\ldots,L$, showing that at each level the conditional distribution $p(\Phi_{\ell-1}|\Psi_{\ell})$ becomes increasingly concentrated, and that this convergence propagates compositionally.
Base case ($\ell=1$): By assumption (1), $\Psi_{1}$ is constructed to minimize the functional:
$\mathcal{L}_{1}=H(\Phi_{0}|\Psi_{1})+\lambda_{1}D_{\mathrm{KL}}\left(\Psi_{1}\,\|\,f_{1}(\Psi_{0}^{(1)},\ldots,\Psi_{0}^{(n_{1})})\right)$
This yields a code $\Psi_{1}$ that provides a more compact representation of $\Phi_{0}$ than any of its inputs $\Psi_{0}^{(i)}$, by assumption (2): $H(\Phi_{0}|\Psi_{1})<H(\Phi_{0}|\Psi_{0}^{(i)}),\quad\forall i$ Furthermore, assumption (3) guarantees bounded representational drift: $D_{\mathrm{KL}}\left(p(\Phi_{0}|\Psi_{1})\,\|\,p(\Phi_{0}|\Psi_{0}^{(i)})\right)\leq\epsilon_{1}$ Together, these imply that the representation $p(\Phi_{0}|\Psi_{1})$ becomes more peaked, while remaining close to its inputs.
Inductive step: Assume that for some level $\ell\geq 1$, the posterior $p(\Phi_{\ell-1}|\Psi_{\ell})$ is sharply concentrated, i.e., has low entropy and small variance: $H(\Phi_{\ell-1}|\Psi_{\ell})\ll H(\Phi_{\ell-1})$ We want to show that this property propagates to the next level $\ell+1$.
From assumption (1), the construction of $\Psi_{\ell+1}$ is again given by:
$\Psi_{\ell+1}=\arg\min_{\Psi}\left[H(\Phi_{\ell}|\Psi)+\lambda_{\ell+1}D_{\mathrm{KL}}\left(\Psi\,\|\,f_{\ell+1}(\Psi_{\ell}^{(1)},\ldots,\Psi_{\ell}^{(n_{\ell+1})})\right)\right]$
Since $\Phi_{\ell}$ is itself a structured summary of $\Phi_{\ell-1}$, the improved compression at level $\ell$ is inherited in the entropy term $H(\Phi_{\ell}|\Psi_{\ell+1})$, and assumption (2) again ensures: $H(\Phi_{\ell}|\Psi_{\ell+1})<H(\Phi_{\ell}|\Psi_{\ell}^{(i)}),\quad\forall i$. Meanwhile, by assumption (3), we have: $D_{\mathrm{KL}}\left(p(\Phi_{\ell}|\Psi_{\ell+1})\,\|\,p(\Phi_{\ell}|\Psi_{\ell}^{(i)})\right)\leq\epsilon_{\ell+1}$, which ensures that the newly formed $\Psi_{\ell+1}$ does not deviate too far in information geometry from its children. This bounded KL divergence, together with the entropy decrease, implies that the posterior distributions $p(\Phi_{\ell}|\Psi_{\ell+1})$ become increasingly sharp and consistent across levels.
Limit behavior: As $\ell\to L$, the repeated application of entropy-minimizing composition with bounded divergence implies that the sequence of posteriors $\{p(\Phi_{\ell-1}|\Psi_{\ell})\}_{\ell=1}^{L}$ forms a convergent sequence in the space of distributions (e.g., under total variation or Wasserstein metric). Since entropy is strictly decreasing and bounded below (by zero), and each KL divergence is finite and contractive, the limit distribution must converge to a delta-like measure concentrated around a low-entropy attractor manifold $\mathcal{M}^{\ast}\subset\mathcal{S}_{L}$, i.e., $\lim_{\ell\to L}p(\Phi_{\ell-1}|\Psi_{\ell})\to\delta_{\mathcal{M}^{\ast}}(\Phi_{\ell-1})$
Conclusion: This proves that hierarchical composition, when performed under entropy reduction and bounded transport divergence, yields a top-level code $\Psi_{L}$ that supports cycle-consistent, context-specific inference via a delta-concentrated representation. This completes the proof. ∎
Appendix H Proof of Theorem 8
Proof.
We proceed by induction over time $t$ and layer $\ell$, showing that the sequence $\{\Phi_{\ell}^{(t)}\}$ converges to a delta-like attractor $\Phi_{\ell}^{\ast}$, such that the conditional distribution $p(\Psi_{\ell}|\Phi_{\ell}^{(t)})$ becomes increasingly peaked and stable.
Step 1: Monotonic entropy contraction at each layer. By assumption (2), we have: $H(\Psi_{\ell}^{(t)}|\Phi_{\ell}^{(t)})>H(\Psi_{\ell}^{(t+1)}|\Phi_{\ell}^{(t+1)})\geq 0$. This implies that the mutual information $I(\Phi_{\ell}^{(t)};\Psi_{\ell}^{(t)})$ increases over time. Since entropy is bounded below by zero, the sequence $\{H(\Psi_{\ell}^{(t)}|\Phi_{\ell}^{(t)})\}$ converges. Therefore, by monotone convergence, the sequence of conditional distributions $\{p(\Psi_{\ell}|\Phi_{\ell}^{(t)})\}$ becomes increasingly concentrated: $\lim_{t\to\infty}p(\Psi_{\ell}|\Phi_{\ell}^{(t)})\to\delta(\Psi_{\ell}-\hat{\Psi}_{\ell}^{\ast})$ for some context prototype $\hat{\Psi}_{\ell}^{\ast}\in\mathbb{R}^{D}$.
Step 2: Contractive inference dynamics. The update rule for content is given by:
$\Phi_{\ell}^{(t+1)}=\mathcal{F}_{\ell}\left(\Phi_{\ell}^{(t)},\hat{\Psi}_{\ell}^{(t)}\right),\quad\hat{\Psi}_{\ell}^{(t)}\sim p(\Psi_{\ell}|Z_{\ell}),\quad Z_{\ell}\sim p(Z_{\ell}|\Phi_{\ell}^{(t)})$
We assume that $\mathcal{F}_{\ell}$ is a stable (e.g., Lipschitz-continuous or contractive) update operator. Since $\hat{\Psi}_{\ell}^{(t)}$ becomes increasingly deterministic as $t\to\infty$, the update dynamics for $\Phi_{\ell}^{(t)}$ converges to a fixed point: $\Phi_{\ell}^{\ast}=\mathcal{F}_{\ell}\left(\Phi_{\ell}^{\ast},\hat{\Psi}_{\ell}^{\ast}\right)$, which implies that $\Phi_{\ell}^{\ast}$ is a stable attractor of the content update process at layer $\ell$.
Step 3: Inductive propagation across layers. From assumption (3), each content vector at layer $\ell$ is computed by a differentiable composition of $n_{\ell}$ content vectors from layer $\ell-1$: $\Phi_{\ell}^{(t)}=f_{\ell}\left(\Phi_{\ell-1}^{(t,1)},\dots,\Phi_{\ell-1}^{(t,n_{\ell})}\right)$. Assuming convergence at layer $\ell-1$, i.e., $\Phi_{\ell-1}^{(t,i)}\to\Phi_{\ell-1}^{\ast(i)}$, and continuity of $f_{\ell}$, it follows that: $\Phi_{\ell}^{(t)}\to\Phi_{\ell}^{\ast}:=f_{\ell}\left(\Phi_{\ell-1}^{\ast(1)},\dots,\Phi_{\ell-1}^{\ast(n_{\ell})}\right)$. Thus, convergence of content representations propagates upward in the hierarchy, preserving delta convergence at each level.
Conclusion: By combining entropy contraction, stable inference dynamics, and layerwise compositionality, we conclude that each $\Phi_{\ell}^{(t)}$ converges to a fixed point $\Phi_{\ell}^{\ast}$ satisfying:
$\Phi_{\ell}^{\ast}=\mathcal{F}_{\ell}(\Phi_{\ell}^{\ast},\hat{\Psi}_{\ell}^{\ast}),\quad p(\Psi_{\ell}|\Phi_{\ell}^{\ast})\approx\delta(\Psi_{\ell}-\hat{\Psi}_{\ell}^{\ast})$
Hence, the full hierarchy converges to a tower of delta-like attractors $\{\Phi_{\ell}^{\ast}\}_{\ell=1}^{L}$, each representing a generalization-optimal scaffold for context prediction. This completes the proof. ∎
Appendix I Proof of Theorem 9
Proof.
We show that the given optimization problem admits a stable solution where all agents converge to a shared latent code $Z\in\mathcal{L}$ that minimizes both the compression cost and the semantic transport cost, while enabling accurate reconstruction of the speaker’s latent state.
Step 1: Compression via Information Bottleneck. Each agent $i$ compresses its internal state $\Phi_{i}\in\mathcal{S}$ into a latent code $Z$ by minimizing the mutual information $I(\Phi_{i};Z)$, balancing expressivity and compactness: $\min_{p(Z|\Phi_{i})}I(\Phi_{i};Z)$. This encourages each agent to produce a compact symbolic representation of its own state that still preserves information relevant for others’ inference.
Step 2: Semantic Transport Cost. Each emitted signal $\Psi_{i\to j}\sim p(\Psi|Z)$ must be interpretable by the receiving agent $j$, which reconstructs $\hat{Z}_{j}\to\hat{\Phi}_{j}$. The mismatch between emitted and received signals is penalized by an optimal transport cost: $D_{\mathrm{OT}}\left(p(\Psi_{i\to j}|Z),\,p(\Psi_{i\to j}|\hat{Z}_{j})\right)$. This term enforces alignment between the speaker’s generated signal and the listener’s interpretation, encouraging convergence of agent-specific decoding functions toward a shared distribution.
Step 3: Social Bootstrapping via Cycle-Consistent Inference. The listener reconstructs $\hat{\Phi}_{j}$ by inferring $\hat{Z}_{j}\to\hat{\Phi}_{j}$. The constraint: $\mathbb{E}_{i\neq j}\left[I(Z;\hat{\Phi}_{j})\right]\geq\mathcal{I}_{0}$ ensures that the reconstructed content $\hat{\Phi}_{j}$ retains a minimum amount of relevant information about the original latent state $\Phi_{i}$. This enforces a global cycle-consistency: each communication loop $\Phi_{i}\to Z\to\Psi\to\hat{Z}_{j}\to\hat{\Phi}_{j}$ must close with minimal semantic loss.
Step 4: Emergence of a Shared Latent Codebook. The optimization is over both the encoder distributions $p(Z|\Phi_{i})$ and signal mappings $p(\Psi_{i\to j}|Z)$. Since all agent pairs $(i,j)$ are jointly minimizing the same global objective and subject to the same reconstruction constraint, the system has a unique pressure: to converge toward a common, shared latent codebook $Z\in\mathcal{L}$ that satisfies all pairwise constraints simultaneously. This shared codebook minimizes redundancy in communication, reduces transport costs, and maximizes inference reliability across the population.
Step 5: Stability under Social Bootstrapping. As agents interact repeatedly, they receive communicative feedback (implicit or explicit) that reinforces mappings between internal states, latent codes, and signals. This bootstrapping process acts as an implicit gradient descent over the shared objective, aligning individual mappings toward a population-level fixed point. By properties of variational convergence in bounded entropy and transport-cost landscapes, and the convexity of mutual information in exponential families, this joint optimization admits a stable equilibrium:
$\{p(Z|\Phi_{i})\}_{i}\to p^{\ast}(Z),\quad\{p(\Psi_{i\to j}|Z)\}_{i,j}\to p^{\ast}(\Psi|Z),\quad\hat{\Phi}_{j}\approx\Phi_{i}$
Thus, agents converge to a cycle-consistent, socially grounded latent codebook $\mathcal{L}$, enabling mutual understanding and inference across the system.
Appendix J Proof of Theorem 10
Proof.
We aim to show that the iterative update
$z_{i}^{(t+1)}=\arg\min_{z_{i}\in\mathcal{L}_{i}}\left[H(\Phi|z_{i})+\lambda\,D_{\mathrm{OT}}\left(p(\Phi|z_{i}),\,p(\Phi|z_{i}^{(t)})\right)\right]$
leads to delta-convergent symbolic assignments for each slot $z_{i}$, and that the full code $Z=(z_{1},\dots,z_{k})$ converges to a fixed compositional representation of $\Phi$.
Step 1: Finite minimization over discrete codes. Since each $\mathcal{L}_{i}$ is a finite discrete set (Assumption 2), and the update selects the minimizer of a continuous objective over a finite set, the update is well-defined and converges in finite time unless oscillation occurs. However, the update is regularized by a transport alignment term, which ensures smooth convergence by penalizing large shifts:
$D_{\mathrm{OT}}(p(\Phi|z_{i}),\,p(\Phi|z_{i}^{(t)}))\text{ small}\Rightarrow z_{i}^{(t+1)}\text{ near }z_{i}^{(t)}$
Step 2: Entropy minimization and contractive update. Each update reduces the conditional entropy: $H(\Phi|z_{i}^{(t+1)})\leq H(\Phi|z_{i}^{(t)})$ unless the current code is already optimal. Since entropy is bounded below (by 0) and decreases over updates, the sequence $\{H(\Phi|z_{i}^{(t)})\}$ is monotonic and convergent. Moreover, the KL-divergence-based contractivity condition (Assumption 1) implies that updates cannot diverge or oscillate indefinitely:
$D_{\mathrm{KL}}(p(\Phi|z_{i}^{(t+1)}),\,p(\Phi|z_{i}^{(t)}))\leq\beta\,D_{\mathrm{KL}}(p(\Phi|z_{i}^{(t)}),\,p(\Phi|z_{i}^{(t-1)}))\quad\text{for some }\beta<1$
This guarantees exponential convergence in distribution.
Step 3: Factorization of the meaning space. By Assumption 3, the joint meaning distribution factorizes: $p(\Phi|Z)\approx\prod_{i=1}^{k}p(\phi_{i}|z_{i})$. Therefore, optimizing each slot $z_{i}$ independently minimizes the full objective:
$\sum_{i}\left[H(\phi_{i}|z_{i})+\lambda\,D_{\mathrm{OT}}(p(\phi_{i}|z_{i}),\,p(\phi_{i}|z_{i}^{(t)}))\right]$
This ensures that convergence in each dimension is sufficient for full convergence of $Z\to\Phi$.
Step 4: Cycle-consistent inference. Assumption 4 enforces communicative feedback: $\Phi\to Z\to\hat{\Phi}\approx\Phi$. This cycle-consistency ensures that symbolic updates that lead to misalignment with the original latent meaning $\Phi$ are penalized and corrected over time. That is, the learning dynamic encourages symbolic consistency that stabilizes meaning.
Step 5: Delta convergence. As $t\to\infty$, the regularized entropy + transport objective is minimized. Since the objective is minimized over discrete sets $\mathcal{L}_{i}$, each sequence $\{z_{i}^{(t)}\}$ converges to a fixed symbol $z_{i}^{\ast}\in\mathcal{L}_{i}$. The limiting distribution becomes:
$p(z_{i}^{(t)})\to\delta(z_{i}^{\ast}),\quad\text{and}\quad p(\Phi|z_{i}^{(t)})\to p(\Phi|z_{i}^{\ast})$
Hence, the full compositional code converges to:
$Z^{\ast}=(z_{1}^{\ast},\dots,z_{k}^{\ast}),\quad\text{with}\quad p(Z^{(t)})\to\delta(Z^{\ast})$
and the resulting codebook minimizes the joint cost: $\sum_{i}\left[H(\phi_{i}|z_{i})+\lambda D_{\mathrm{OT}}(p(\phi_{i}|z_{i}),\,p(\phi_{i}|z_{i}^{(t)}))\right]$.
Conclusion: Therefore, under finite support, entropy contraction, contractive updates, and factorized latent semantics, symbolic codes converge to delta-like attractors encoding distinct semantic factors. This yields a stable, interpretable, compositional representation of structured meaning. ∎

### v2

@@
 Within this framework, cycle formation resolves the Information Bottleneck in Optimal Transport (iBOT) by coordinating bottom-up contextual disambiguation with top-down content reconstruction.
-This dual process is implemented via a Rao-Blackwellized inference strategy, formalized as variational entropy minimization, which steers the system toward minimal joint uncertainty while preserving the directionality of inference.
-Crucially, local cycle completion gives rise to temporally extended chain formation (temporal bootstrapping) and spatially distributed hierarchical composition (spatial bootstrapping), enabling the system to simulate goals, evaluate counterfactuals, and iteratively refine internal models.
+This dual process is implemented via a memory-amortized inference (MAI) strategy, formalized as variational entropy minimization, which steers the system toward minimal joint uncertainty while preserving the directionality of inference.
+Importantly, local cycle completion gives rise to temporally extended chain formation (temporal bootstrapping) and spatially distributed hierarchical composition (spatial bootstrapping), enabling the system to simulate goals, evaluate counterfactuals, and iteratively refine internal models.
 To formalize this bootstrapped memory process, we prove a Delta Convergence Theorem showing that recursive entropy minimization over time yields structured memory states that converge to delta-like distributions over latent manifolds.
@@
 As such, inverted inference provides a powerful mechanism for resolving iBOT by avoiding the compression of entangled context and instead organizing cognition around reusable, generalizable seeds of content.
-Inverted inference gains its full expressive power when extended through recursive bootstrapping across both temporal and spatial hierarchies li2025Inverted .
+Inverted inference gains its full expressive power when extended through recursive bootstrapping across temporal and spatial hierarchies li2025Inverted .
 In the temporal domain, each content state is updated not in isolation but as part of a dynamic sequence, where structure learned at time $t$ serves as the scaffold for inference at $t+1$.
@@
 Here, the structure is built recursively by assembling local regularities into increasingly abstract representations, satisfying both inference sufficiency and compositional coherence.
-The unification of these two bootstrapping processes, temporal and spatial, forms a multiscale inference engine wherein each layer constrains the next, and structure is not only inferred but grown, recursively and contextually, in space and time.
+The unification of these two bootstrapping processes forms a multiscale inference engine wherein each layer constrains the next, and structure is not only inferred but grown, recursively and contextually, in space and time.
 Recursive bootstrapping induces a natural convergence dynamic between content and context representations.
@@
 Lemma 2 (Entropy Decomposition Lemma with Variational Formulation).
-[…]
+li2025CCUP Let $\Psi$ denote a high-entropy contextual variable and $\Phi$ denote a structured, low-entropy content variable such that their joint distribution $p(\Phi,\Psi)$ is well-defined.
+Then, for any inference system that seeks to minimize the uncertainty over $\Phi$ and $\Psi$, the optimal inference policy proceeds by: 1) Minimizing the joint entropy via decomposition: $H(\Phi,\Psi)=H(\Phi)+H(\Psi|\Phi)$; 2) Under the assumption $H(\Phi)\ll H(\Psi)$, the conditional entropy $H(\Psi|\Phi)$ dominates and becomes the principal target of minimization; 3) This induces a directional alignment: inference should proceed by first establishing $\Phi$, then using it to constrain $\Psi$, such that: $\Phi\leadsto\Psi\quad\text{(structure precedes specificity)}$.
+Moreover, the variational inference objective: $\mathcal{F}[q]=\mathbb{E}_{q(Z|\Psi)}[-\log p(\Psi|Z)]+D_{\mathrm{KL}}(q(Z|\Psi)\|p(Z|\Phi))$ is minimized when the prior $p(Z|\Phi)$ encodes low-entropy structure aligned with $\Psi$.
 How IB in OT Arises from the Broken Symmetry.
@@
 li2025Inverted Let $\Psi$ denote a high-entropy context variable and $\Phi$ a low-entropy content variable, with bidirectional inference maps $q(\Phi|\Psi)$ (inference) and $p(\Psi|\Phi)$ (generation).
-The pair $(\Psi,\Phi)$ is said to be aligned if the following equivalent conditions hold: 1) Cycle Consistency: $D_{\mathrm{KL}}(\Psi\,\|\,\hat{\Psi})\leq\varepsilon,D_{\mathrm{KL}}(\Phi\,\|\,\hat{\Phi})\leq\varepsilon$, where $\hat{\Phi}=q(\Phi|\Psi)$ and $\hat{\Psi}=p(\Psi|\hat{\Phi})$, and $\varepsilon\geq 0$ is a small alignment tolerance.
+The pair $(\Psi,\Phi)$ is said to be aligned if the following equivalent conditions hold: 1) Cycle Consistency: $D_{\mathrm{KL}}(\Psi\,\|\,\hat{\Psi})\leq\varepsilon,D_{\mathrm{KL}}(\Phi\,\|\,\hat{\Phi})\leq\varepsilon$, where $\hat{\Phi}=q(\Phi|\Psi)$ $\hat{\Psi}=p(\Psi|\hat{\Phi}), and$ $\varepsilon\geq 0$ is a small alignment tolerance.
 2) Mutual Information Saturation: $I(\Psi;\Phi)\approx H(\Phi)\approx H(\Psi)$, indicating that context and content share nearly all their informational degrees of freedom.
@@
 Note that a shortcut is generated by predictive coding, which implements the updating of content without completing the cycle.
-[…]
+To make inverted inference tractable and expressive, we introduce a latent variable $Z$ that serves as a variational mediator between structured content $\Phi$ and ambiguous context $\Psi$ (please refer to Fig.
+1).
+Rather than modeling the direct mapping $p(\Psi|\Phi)$, which may be highly multi-modal or entangled, we define a two-stage generative model: $p(\Psi|\Phi)=\int p(\Psi|Z)\,p(Z|\Phi)\,dZ$, where $p(Z|\Phi)$ samples generalizable latent codes conditioned on compact structure and $p(\Psi|Z)$ decodes those codes into plausible contexts.
+The latent variable $Z$ functions as an information-theoretic stepping stool that facilitates the flow from abstract content $\Phi$ to entangled context $\Psi$.
+By mediating this transformation through an intermediate space, $Z$ enables a structured detour that preserves latent compositionality while overcoming the direct compression bottleneck imposed by high-dimensional context.
+This factorization decomposes the inference problem into two manageable steps: generalization via $p(Z|\Phi)$ and contextualization via $p(\Psi|Z)$, making efficient and precise structure-to-context alignment possible.
 More rigorously, variational formulation introduces a recognition model $q(Z|\Psi)$ and yields an evidence lower bound (ELBO) for learning: $\log p(\Psi|\Phi)\geq\mathbb{E}_{q(Z|\Psi)}[\log p(\Psi|Z)]-D_{\mathrm{KL}}(q(Z|\Psi)\,\|\,p(Z|\Phi))$.
@@
 Lemma 4 (Predictive Coding as a Half-Cycle Right Refinement).
-Let $\Phi\in\mathcal{S}$ denote structured content, and let $\Psi\in\mathcal{C}$ be contextual observations.
-Consider the latent variable $Z\in\mathcal{Z}$ mediating the generative path from structure to context: $p(Z|\Phi),\quad p(\Psi|Z)$.
-Define the forward generative prediction: $\hat{\Psi}=\mathbb{E}_{Z\sim p(Z|\Phi)}\left[\mathbb{E}_{\Psi\sim p(\Psi|Z)}[\Psi]\right]$, and define the backward recognition refinement via amortized inference: $\hat{Z}=q(Z|\Psi),\quad\Phi^{\prime}=q(\Phi|\hat{Z})$.
-Assume the update $\Phi\to\Phi^{\prime}$ minimizes a context-aligned prediction error: $\Phi^{\prime}=\arg\min_{\Phi}D(\Psi\,\|\,\hat{\Psi}(\Phi))$, where $D$ is a divergence metric (e.g., KL or squared error).
-Then the half-cycle operator: $\Phi\xrightarrow{p(Z|\Phi)\rightarrow p(\Psi|Z)}\hat{\Psi}\xrightarrow{q(Z|\hat{\Psi})\rightarrow q(\Phi^{\prime}|Z)}\Phi^{\prime}$ is equivalent to a predictive coding update that refines $\Phi$ by descending the gradient of prediction error on $\Psi$.
+[…]
 The preceding lemma formalizes predictive coding as a half-cycle right refinement process, in which structured content $\Phi$ is updated through a generative–recognition loop grounded in prediction error minimization.
@@
 Theorem 1 (Breaking the Curse of Dimensionality via Half-Cycle Right Inference).
-Let $\Psi\in\mathbb{R}^{D}$ denote a high-dimensional context variable, and let $\Phi\in\mathbb{R}^{d}$ denote a structured latent representation with $d\ll D$.
-Define a latent bottleneck variable $Z\in\mathbb{R}^{k}$ with $d\leq k\ll D$, and consider a cycle-consistent inference scheme: $\Phi^{(t)}\xrightarrow{p(Z|\Phi)}Z^{(t)}\xrightarrow{p(\Psi|Z)}\hat{\Psi}^{(t)}\xrightarrow{q(Z|\hat{\Psi})}\hat{Z}^{(t)}\xrightarrow{q(\Phi|\hat{Z})}\Phi^{(t+1)}$.
-Assume 1) The joint model $p(\Psi,Z,\Phi)$ factorizes through low-dimensional conditional distributions; 2) The update $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$ is entropy-decreasing: $H(\Psi|\Phi^{(t+1)})<H(\Psi|\Phi^{(t)})$; 3) The reconstruction steps $q(Z|\hat{\Psi})$, $q(\Phi|\hat{Z})$ are amortized and Lipschitz-continuous, then the inference process converges to a structured representation $\Phi^{\ast}$ that achieves predictive sufficiency: $H(\Psi|\Phi^{\ast})\approx\min_{\Phi}H(\Psi|\Phi)$ by avoiding direct marginalization over $\Psi\in\mathbb{R}^{D}$ and reducing inference complexity from $\mathcal{O}(\exp(D))$ to $\mathcal{O}(k^{2})$ under structured priors and amortization.
+[…]
 As an application of Theorem 1, we use two classical examples in spatial and visual inference to illustrate how half-cycle inference seeded from goal structure can break the curse of dimensionality by replacing forward trajectory expansion with predictive refinement guided by internal priors.
@@
 Let $V_{t}\in\mathcal{X}$ be an evolving representation at iteration $t$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ is a compact, convex space.
-Suppose the update rule takes the bootstrapped form: $V_{t+1}=\mathcal{F}(V_{t},\widetilde{E}(V_{t}))$, where $\mathcal{F}:\mathcal{X}\times\mathcal{Y}\to\mathcal{X}$ is a deterministic update operator, $\widetilde{E}:\mathcal{X}\to\mathcal{Y}$ is a surrogate or predictive operator and $\mathcal{F}$ is contractive in its first argument: there exists $10<\gamma<$ 1 such that $\|\mathcal{F}(x_{1},y)-\mathcal{F}(x_{2},y)\|\leq\gamma\|x_{1}-x_{2}\|\quad\text{for all }x_{1},x_{2}\in\mathcal{X},y\in\mathcal{Y}$.
+Suppose the update rule takes the bootstrapped form: $V_{t+1}=\mathcal{F}(V_{t},\widetilde{E}(V_{t}))$, where $\mathcal{F}:\mathcal{X}\times\mathcal{Y}\to\mathcal{X}$ is a deterministic update operator, $\widetilde{E}:\mathcal{X}\to\mathcal{Y}$ is a surrogate or predictive operator and $\mathcal{F}$ is contractive in its first argument: there exists $0<\gamma<1$ such that $\|\mathcal{F}(x_{1},y)-\mathcal{F}(x_{2},y)\|\leq\gamma\|x_{1}-x_{2}\|\quad\text{for all }x_{1},x_{2}\in\mathcal{X},y\in\mathcal{Y}$.
 Then the sequence $\{V_{t}\}_{t=0}^{\infty}$ converges to a unique fixed point $V^{\ast}\in\mathcal{X}$, satisfying: $V^{\ast}=\mathcal{F}(V^{\ast},\widetilde{E}(V^{\ast}))$.
@@
 Example 4 (Focused Ultrasound as Restoration of Delta Convergence in Parkinson’s Disease).
-[…]
+Let $\Phi^{(t)}\in\mathcal{S}\subset\mathbb{R}^{d}$ denote the structured motor plan at time $t$, updated via recursive inference: $\Phi^{(t+1)}=\arg\min_{\Phi}\left[H(\Psi^{(t)}\mid\Phi)+\lambda\,\mathcal{D}(p(Z\mid\Phi),r(Z))\right]$, where $\Psi^{(t)}$ is the proprioceptive or external motor feedback, $Z$ is a latent motor intention, and $r(Z)$ is a prior over motor programs.
+In PD, due to dopaminergic degeneration and basal ganglia dysfunction, the recursive update operator becomes dysregulated and non-contractive: $\exists\,\Phi_{1},\Phi_{2}\in\mathcal{S},\quad\|\mathcal{F}(\Phi_{1},\hat{\Psi})-\mathcal{F}(\Phi_{2},\hat{\Psi})\|\geq\|\Phi_{1}-\Phi_{2}\|$, causing the sequence $\{\Phi^{(t)}\}$ to diverge or oscillate, leading to motor tremor and instability tarakad2019essential .
+Now, let a therapeutic intervention such as focused ultrasound (FUS) or deep brain stimulation (DBS) perlmutter2006deep modify the neural gain or effective damping of subcortical structures, particularly in the thalamus or subthalamic nucleus.
+This is modeled as restoring contractivity of the update operator $\mathcal{F}_{\mathrm{stim}}$: $\|\mathcal{F}_{\mathrm{stim}}(\Phi_{1},\hat{\Psi})-\mathcal{F}_{\mathrm{stim}}(\Phi_{2},\hat{\Psi})\|\leq\gamma\|\Phi_{1}-\Phi_{2}\|,\quad 0<\gamma<1$.
+Under these modified dynamics, the inference loop now satisfies the conditions for Delta Convergence: $\lim_{t\to\infty}p(Z\mid\Phi^{(t)})=\delta(Z-Z^{\ast})$, leading to stable convergence of the motor plan: $\Phi^{(t)}\to\Phi^{\ast},\quad\text{where }\Phi^{\ast}\text{ is a self-consistent, tremor-free attractor}$.
+From an inferential dynamics perspective, ultrasound and DBS therapies stabilize recursive updates in motor planning by restoring the contractive geometry of latent-to-context inference, allowing PD patients to recover precise, stable motor control.
 4 Spatiotemporal Bootstrapping for Fast and Slow Thinking
@@
 Theorem 7 (Delta Convergence of Hierarchical Composition).
-Let $\{\Psi_{\ell}\}_{\ell=0}^{L}$ be a hierarchy of latent variables forming a compositional structure under the conditions of Theorem LABEL:thm:HCSS.
+Let $\{\Psi_{\ell}\}_{\ell=0}^{L}$ be a hierarchy of latent variables forming a compositional structure under the conditions of Theorem 3.
 Suppose: 1) Each level $\Psi_{\ell}$ is obtained via entropy-aligned composition: $\Psi_{\ell}=\arg\min_{\Psi}\left[H(\Phi_{\ell-1}|\Psi)+\lambda_{\ell}D_{\mathrm{KL}}\left(\Psi\,\|\,f_{\ell}(\Psi_{\ell-1}^{(1)},\ldots,\Psi_{\ell-1}^{(n_{\ell})})\right)\right]$; 2) The conditional entropy decreases at each level: $H(\Phi_{\ell-1}|\Psi_{\ell})<H(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)}),\quad\forall i$; 3) The KL divergence across levels is uniformly bounded: $D_{\mathrm{KL}}\left(p(\Phi_{\ell-1}|\Psi_{\ell})\,\|\,p(\Phi_{\ell-1}|\Psi_{\ell-1}^{(i)})\right)\leq\epsilon_{\ell}$, then, in the limit $\ell\to L$, the top-level representation $\Psi_{L}$ converges in distribution to a delta-like measure concentrated around a minimal uncertainty manifold $\mathcal{M}^{\ast}\subset\mathcal{S}_{L}$, such that: $\lim_{\ell\to L}p(\Phi_{\ell-1}|\Psi_{\ell})\to\delta_{\mathcal{M}^{\ast}}(\Phi_{\ell-1})$ This implies that inference becomes maximally specific with respect to context, while remaining globally structured and compositionally abstract.
@@
 Example 6 (Memory Restoration as Recovery of Cycle-Consistent Inference).
-[…]
+Let $\Phi\in\mathcal{S}\subset\mathbb{R}^{d}$ represent structured latent memory (e.g., semantic or episodic code), $Z\in\mathbb{R}^{k}$ an intermediate latent variable, and $\Psi\in\mathcal{C}\subset\mathbb{R}^{D}$ denote high-dimensional contextual observations.
+In the C2HM framework, memory is modeled as a recursive inference cycle: $\Phi^{(t)}\xrightarrow{p(Z|\Phi)}Z^{(t)}\xrightarrow{p(\Psi|Z)}\hat{\Psi}^{(t)}\rightarrow Z^{(t)\prime}\xrightarrow{q(\Phi|Z^{(t)\prime})}\Phi^{(t+1)}$, where the update $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$ refines structured memory via simulation and correction.
+In Alzheimer’s disease (AD) scheltens2021alzheimer , this cycle is disrupted by multiple interacting mechanisms: 1) Accumulation of $\beta$-amyloid and tau impairs the structural integrity of $\mathcal{F}$, reducing plasticity and recursive refinement; 2) Network disconnection decouples the mapping $p(Z|\Phi)$, weakening simulation and pattern completion; 3) Chronic inflammation introduces stochastic noise in $p(\Psi|Z)$, increasing entropy in context predictions.
+Effective memory restoration, therefore, must target not just structural damage, but the recovery of the inference loop palop2016network .
+Let $\mathcal{T}=\{\text{Plasticity},\text{Stimulation},\text{Anti-Tau},\text{Simulation Training}\}$ denote a set of interventions.
+Each therapy $\tau\in\mathcal{T}$ acts to restore one or more components of the bootstrapping cycle.
+When these components are restored, the entropy of the forward prediction $H(\Psi|\Phi^{(t)})$ and the reconstruction divergence $D(\Phi^{(t+1)}\|\Phi^{(t)})$ both decrease over cycles.
+The system converges to a fixed point $\Phi^{\ast}$ such that: $\Phi^{\ast}=\arg\min_{\Phi}\mathbb{E}_{Z\sim p(Z|\Phi)}\left[-\log p(\Psi|Z)\right]+\lambda\cdot D_{\mathrm{KL}}(p(Z|\Phi)\|r(Z))$, where $r(Z)$ is a reference prior enforcing compactness.
+Therefore, memory restoration is best understood not as retrieval of stored data, but as recovery of an active, entropy-minimizing inference process.
 5 Social Bootstrapping and the Origin of Language
-From Spatiotemporal Bootstrapping to Social Bootstrapping.
+5.1 From Spatiotemporal Bootstrapping to Social Bootstrapping
 While spatiotemporal bootstrapping enables individual agents to construct coherent internal models by integrating information across scales of space (e.g., object composition) and time (e.g., episodic memory), such inference remains fundamentally bounded by the agent’s own sensory history and computational constraints.
@@
 Let $\hat{\Phi}\in\mathcal{S}$ denote the listener’s inferred internal state based on observing $\Psi$.
-The Information Bottleneck for Social Communication is the problem of finding a stochastic encoding $p(\Psi|\Phi)$ that minimizes the mutual information between the internal state and the signal, $\min_{p(\Psi|\Phi)}I(\Phi;\Psi)$, subject to a constraint on the relevance of the signal to the listener’s inference task: subject to $\text{subject to }I(\Psi;\hat{\Phi})\geq\mathcal{I}_{0}$, where $\mathcal{I}_{0}$ is a lower bound on the amount of relevant information that must be preserved for successful communication.
+The Information Bottleneck for Social Communication is the problem of finding a stochastic encoding $p(\Psi|\Phi)$ that minimizes the mutual information between the internal state and the signal, $\min_{p(\Psi|\Phi)}I(\Phi;\Psi)$, subject to a constraint on the relevance of the signal to the listener’s inference task: $\text{subject to }I(\Psi;\hat{\Phi})\geq\mathcal{I}_{0}$, where $\mathcal{I}_{0}$ is a lower bound on the amount of relevant information that must be preserved for successful communication.
 This tradeoff captures the core challenge of social communication: compressing internal content into efficient signals that remain maximally informative under the listener’s reconstruction process.
@@
 Let $Z\in\mathcal{L}$ denote a shared latent linguistic representation.
-Language acts as an intermediating variable $Z$ that solves the Information Bottleneck (IB) for Optimal Transport (OT) in social communication, by satisfying the following: 1) Compression: The map $\Phi\to Z$ minimizes the encoding cost while preserving communicative relevance: $\min_{p(Z|\Phi)}I(\Phi;Z)\quad\text{subject to}\quad I(Z;\hat{\Phi})\geq\mathcal{I}_{0}$; 2) Transport: The signal $\Psi$ is generated from $Z$ in a way that minimizes semantic transport mismatch: $\min_{p(\Psi|Z)}D_{\mathrm{OT}}\big{(}p(\Psi|Z),\,p(\Psi|\hat{Z})\big{)}$; 3) Cycle-consistency: The listener infers $\hat{Z}\to\hat{\Phi}$, and successful communication occurs when $\hat{\Phi}\approx\Phi$, closing the inference loop.
+Language acts as an intermediating variable $Z$ that solves the Information Bottleneck (IB) for Optimal Transport (OT) in social communication, by satisfying the following: 1) Compression: The map $\Phi\to Z$ minimizes the encoding cost while preserving communicative relevance: $\min_{p(Z|\Phi)}I(\Phi;Z)\quad\text{subject to}\quad I(Z;\hat{\Phi})\geq\mathcal{I}_{0}$; 2) Transport: The signal $\Psi$ is generated from $Z$ in a way that minimizes semantic transport mismatch: $)\min_{p(\Psi|Z)}D_{\mathrm{OT}}\big{(}p(\Psi|Z),\,p(\Psi|\hat{Z}$$\big{)}$; 3) Cycle-consistency: The listener infers $\hat{Z}\to\hat{\Phi}$, and successful communication occurs when $\hat{\Phi}\approx\Phi$, closing the inference loop.
 Hence, language defines a shared latent space $\mathcal{L}$ that compresses, transports, and reconstructs structured meaning across agents.
@@
 (a) Single-Agent IB–OT Triangle
-$\Phi_{1}$$\Phi_{2}$$\Phi_{3}$hared Latent Language $Z\in\mathcal{L}$$\Psi_{1\to 2}$$\Psi_{2\to 3}$$\Psi_{3\to 1}$ulti-Agent Alignment via Language
+$\Phi_{1}$$\Phi_{2}$$\Phi_{3}$Shared Latent Language $Z\in\mathcal{L}$$\Psi_{1\to 2}$$\Psi_{2\to 3}$$\Psi_{3\to 1}$Multi-Agent Alignment via Language
 (b) Social Bootstrapping Across Agents
@@
 Let $Z=(z_{1},z_{2},z_{3})\in\mathcal{L}_{1}\times\mathcal{L}_{2}\times\mathcal{L}_{3}$ be a compositional symbolic representation of $\Phi$, where each $z_{i}$ is selected from a finite vocabulary slot corresponding to one semantic factor.
-Across languages, different mappings from $\Phi\to Z\to\Psi$ emerge, but each converges to a delta-like attractor within its symbolic system: 1) English (Satellite-framed): $z_{1}=\texttt{run}$, $z_{2}=\texttt{into}$, $z_{3}=\texttt{the house}$Rightarrow “The girl ran into the house.”
+Across languages, different mappings from $\Phi\to Z\to\Psi$ emerge, but each converges to a delta-like attractor within its symbolic system: 1) English (Satellite-framed): $z_{1}=\texttt{run}$, $z_{2}=\texttt{into}$, $z_{3}=\texttt{the house}$ $\Rightarrow$ “The girl ran into the house.”
 […]
-[…]
+3) Tzeltal (Mayan): Path is encoded in the verb root, manner added via modifier, goal via spatial particle: $z_{1}=\texttt{bajt'}$ (go.in), $z_{2}=\texttt{lok'el}$ (running), $z_{3}=\texttt{ta na}$ (to house) $\Rightarrow$ Compositional sentence: “She ran into the house.”
 Despite these typological differences, each language exhibits: 1) Factorization: a mapping from latent semantics $\Phi$ to symbolic slots $Z$; 2) Stability: repeated use leads to convergence to discrete symbol choices $z_{i}^{\ast}$; 3) Expressive equivalence: the capacity to communicate the same structured meaning across systems.
 This supports the conclusion of Theorem 10: the symbolic code $Z$ in each language converges to a fixed point minimizing joint entropy and transport distortion over semantic structure.
-6 Conclusion: Information Must Flow
+5.2 Chomsky Hierarchy as an Inference-Memory Spectrum under CCUP
+Traditionally, the Chomsky hierarchy classifies formal languages by their generative power:
+$\text{Regular}\subset\text{Context-Free}\subset\text{Context-Sensitive}\subset\text{Recursively Enumerable}.$
+From the perspective of the CCUP, this hierarchy reflects a graded scale of cycle complexity, memory demands, and uncertainty minimization within latent inference loops:
+  • Regular Languages: Involve flat, local inference with no nested structure.
+MAI retrieval is nearly stateless, and latent trajectories do not form topological cycles (i.e., $H_{k}=0$ for $k\geq 1$).
+  • Context-Free Languages: Require recursive structure such as balanced parentheses or nested phrases.
+These correspond to persistent 1-cycles $\gamma]\in H_{1}(\mathcal{S})$, allowing structured reuse of latent memory paths.
+  • Context-Sensitive Languages: Depend on both local and global context, requiring intertwined inference paths with conditional dependencies.
+These entail higher-dimensional cycles and joint minimization of $H(\Psi,\Phi)$.
+  • Recursively Enumerable Languages: Represent open-ended inference without guaranteed closure.
+Under CCUP, these correspond to divergent latent trajectories with no convergence in homology—i.e., inference proceeds under unbounded uncertainty.
+Corollary 4 (CCUP Interpretation of the Chomsky Hierarchy).
+Let $\mathcal{S}\subset\mathbb{R}^{d}$ be the latent manifold of inference cycles under CCUP, and let $H_{k}(\mathcal{S})$ denote its persistent homology.
+Then: 1) Regular languages: $H_{k}(\mathcal{S})=0$ for $k\geq 1$; inference proceeds without latent cycles.
+2) Context-free languages: Require 1-cycles $\gamma]\in H_{1}(\mathcal{S})$ for recursive structure reuse.
+3) Context-sensitive languages: Require structured memory trajectories embedded in higher-dimensional cycles $H_{\geq 1}(\mathcal{S})$.
+4) Recursively enumerable languages: Correspond to divergent inference loops with no topological stabilization, representing maximal uncertainty.
+Thus, the Chomsky hierarchy encodes increasing complexity of cycle-consistent inference, memory structure, and entropy minimization under CCUP.
+5.3 Recursion as a Context-Driven Phase Transition under CCUP
+The long-standing debate over the universality of recursion in language, exemplified by Chomsky’s claim that recursion is a universal feature of human language, and Everett’s assertion that the Pirahã language lacks it, can be reframed under the CCUP.
+Rather than treating recursion as a hardwired syntactic necessity, CCUP views recursion as an emergent structural response to contextual uncertainty.
+In this framework, recursive inference loops arise only when the entropy of context $H(\Psi)$ cannot be adequately compressed by flat, non-recursive content representations $\Phi$.
+Pirahã as a Low-Entropy Context Language.
+According to Everett everett2005cultural , Pirahã speakers restrict linguistic expression to immediate, lived experience, avoiding abstract referential structures or complex clause embedding.
+Under CCUP, this corresponds to a low-entropy context distribution $\Psi$, where incoming signals are sufficiently specific that shallow inference suffices.
+Because uncertainty is low, there is no need for recursive inference structures to emerge: the loop between context $\Psi$ and inferred content $\Phi$ quickly converges, and the latent space $\mathcal{S}$ remains locally linear.
+Recursion as Loop Closure in Latent Inference.
+In CCUP, recursion is interpreted as a loop closure in latent space: a re-entry of previously inferred content into the current inference cycle.
+Formally, recursion corresponds to the emergence of persistent 1-cycles $\gamma]\in H_{1}(\mathcal{S})$, where the latent trajectory reuses structural components in a context-sensitive manner.
+This interpretation implies that recursion is not a universal constant, but a context-dependent phase transition in the dynamics of uncertainty minimization.
+Cognitive Universality vs. Linguistic Expression.
+Even if Pirahã lacks overt syntactic recursion, its speakers may still employ recursive inference structures in domains like spatial reasoning or social cognition.
+CCUP separates cognitive recursion, emergent from high-dimensional uncertainty resolution, from linguistic recursion chomsky2014minimalist , its grammatical surface realization.
+The absence of recursion in Pirahã syntax thus reflects a cultural constraint on communicative context, not a limitation in cognitive architecture.
+Corollary 5 (Recursion as a Phase Transition under CCUP).
+Let $\Psi$ denote a distribution over linguistic or perceptual contexts, and $\Phi$ a latent content space inferred under CCUP.
+Define a recursive inference structure as the emergence of a nontrivial cycle $\gamma\in H_{1}(\mathcal{S})$, formed by the repeated reuse of content $\Phi$ in successive context–content inference loops.
+Then we have: 1) If $H(\Psi)\leq H_{\text{flat}}$, where $H_{\text{flat}}$ is the entropy compressible by non-recursive inference, then inference converges to shallow content representations $\Phi\in\mathcal{S}$ with no persistent cycles.
+2) If $H(\Psi)>H_{\text{flat}}$, recursive reuse of latent structures becomes necessary, and inference transitions to loop-forming dynamics that generate persistent homology classes $\gamma]\in H_{1}(\mathcal{S})$.
+3) Therefore, recursion is not a binary universal but a phase transition triggered by excess contextual entropy, i.e., $\text{Recursion Emerges}\iff H(\Psi)>H_{\text{flat}}$.
+This corollary aligns with a graded, structure-aware theory of linguistic universals: recursion is not a genetically hardwired feature, but an emergent strategy for minimizing uncertainty under demanding contextual conditions.
+Languages like Pirahã, which restrict referential complexity, simply operate below the entropy threshold required to induce recursive inference, as shown in Fig.
+3.
+[…]
+Figure 3: Recursion emerges as a phase transition when context entropy $H(\Psi)$ exceeds the compression capacity $H_{\text{flat}}$ of shallow inference.
+6 Conclusions and Perspectives
+Inference Bottleneck of Transport: Information must Flow.
 This paper has advanced a unified framework for understanding the directional flow of information in cognitive and communicative systems, grounded in the interplay between the IB, OT, and the CCUP.
@@
 From the recursive architecture of individual cognition to the symbolic scaffolds of social communication, overcoming iBOT is the central computational challenge that shapes the evolution of intelligent systems.
+Toward Human-Level AGI: A Framework of Necessary Components.
+The emergence of human-level AGI kurzweil2005singularity or superintelligence bostrom2024superintelligence is not a matter of scale alone; it is a matter of how information flows.
+Intelligence arises when information flows are structured, filtered, and recursively aligned across perception, memory, and action.
+At the heart of this process is the resolution of the data bottleneck, which cannot be overcome by brute force.
+It requires architectures, such as the CCUP, that maintain cycle-consistent information flow between structured content and ambiguous context, allowing the system to compress, predict, and generalize from limited data.
+This must be coupled with latent world models hafner2019dream that leverage symbolic mediation, enabling information to flow across compositional levels: from sensory input to abstract reasoning, from present states to counterfactual futures.
+Such models curate compact representations that support hierarchical inference and flexible adaptation.
+Importantly, information flow must be shaped by goal-seeded active simulation, not merely reacting to data, but proactively generating, testing, and refining predictions driven by internal objectives.
+This self-directed simulation must unfold within a spatiotemporal bootstrapping architecture, where short-term learning scaffolds long-term memory, and fast cycles of perception–action integrate over extended horizons to produce coherent behavior.
+Finally, to reach true generality, AGI must align its internal information flows with human semantic goals.
+It must not merely optimize task metrics, but interpret and navigate open-ended environments with respect to human meaning, values, and intent.
+In this view, AGI emerges not as a monolith of compute but as a structured information-processing organism, where architecture, simulation, and alignment coevolve to sustain purposeful, context-aware inference under real-world constraints.
+Rather than treating intelligence as the accumulation of knowledge or optimization power, we frame it as a frictionless, cycle-consistent information flow, abstracted by the Energy–Entropy–Structure–Meaning Cycle, that curates structured meaning from environmental entropy and recursively uses that meaning to direct future inference and control.
 Appendix A Proof of Theorem 1
@@
 Therefore, $\Phi^{(t)}\to\Phi^{\ast}$ in the space of structured representations, and $\Phi^{\ast}$ minimizes conditional entropy.
-Step 3: Convergence Under Lipschitz and Amortization Assumptions.
-Let us define the update operator $\mathcal{F}\colon\mathbb{R}^{d}\times\mathbb{R}^{D}\to\mathbb{R}^{d}$ via: $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)},\hat{\Psi}^{(t)})$, where $\hat{\Psi}^{(t)}\sim p(\Psi|Z^{(t)})$, $Z^{(t)}\sim p(Z|\Phi^{(t)})$.
-Since both $q(Z|\hat{\Psi})$ and $q(\Phi|\hat{Z})$ are assumed to be amortized and Lipschitz-continuous, and the composite operator $\mathcal{F}$ is entropy-decreasing, Banach’s fixed-point theorem ensures that the iterative update converges to a unique fixed point $\Phi^{\ast}$ in $\mathbb{R}^{d}$.
+[…]
 Step 4: Computational Advantage.
@@
 Step 2 (Contractive mapping convergence). Since $\mathcal{F}$ is contractive in the complete metric space $(\mathcal{S},d)$, i.e.,
-$d(\mathcal{F}(\Phi_{1},\Psi),\mathcal{F}(\Phi_{2},\Psi))\leq\lambda d(\Phi_{1},\Phi_{2}),\quad\text{for all }\Phi_{1},\Phi_{2}\in\mathcal{S},\text{ with }0<\lambda<1,$
+[…]
 it follows from the Banach Fixed-Point Theorem that the sequence $\{\Phi^{(t)}\}$ converges to a unique fixed point $\Phi^{*}\in\mathcal{S}$ such that: $\Phi^{*}=\mathcal{F}(\Phi^{*},\Psi^{*})$, where $\Psi^{*}:=\lim_{t\to\infty}\Psi^{(t)}$ (assumed to converge or fluctuate within bounded entropy).
@@
 Furthermore, the update minimizes a Bregman-like proximal step around $\Psi^{(t)}$, which satisfies: $D_{\mathrm{KL}}(\Psi^{\ast}\,\|\,\Psi^{(t+1)})<D_{\mathrm{KL}}(\Psi^{\ast}\,\|\,\Psi^{(t)})$ whenever $\Psi^{(t)}\neq\Psi^{\ast}$, due to strict convexity and the entropy gap.
-[…]
+Step 3: Descent and Convergence.
+This implies $\{\mathcal{D}^{(t)}\}$ is a strictly decreasing, bounded-below sequence.
+Hence, $\mathcal{D}^{(t)}\to 0$, i.e., $\lim_{t\to\infty}D_{\mathrm{KL}}(\Psi^{\ast}\,\|\,\Psi^{(t)})=0$, which implies $\Psi^{(t)}\xrightarrow{\text{weak-*}}\Psi^{\ast}$.
 Step 4: Collapse to Delta Measure.
@@
 Proof.
-[…]
+Define the loss function: $\mathcal{L}(\Phi):=H(\Psi|\Phi)+\lambda D_{\mathrm{OT}}(p(\Psi|\Phi),\nu)$ By assumptions (1) and (2), both terms are convex and continuous, so $\mathcal{L}(\Phi)$ is convex and continuous on the compact domain $\mathcal{S}$.
+Hence, the minimizer $\Phi^{*}=\arg\min_{\Phi\in\mathcal{S}}\mathcal{L}(\Phi)$ exists.
+If either term is strictly convex, then $\Phi^{*}$ is unique.
+Define the deterministic update operator $\mathcal{F}:\mathcal{S}\rightarrow\mathcal{S}$ by: $\mathcal{F}(\Phi^{(t)})=\arg\min_{\Phi\in\mathcal{S}}\mathcal{L}(\Phi)$ Then the update rule becomes: $\Phi^{(t+1)}=\mathcal{F}(\Phi^{(t)})$ This defines a best-response dynamic that strictly decreases $\mathcal{L}(\Phi^{(t)})$, unless $\Phi^{(t)}=\Phi^{*}$.
+Since $\mathcal{L}$ is lower-bounded (as both entropy and transport cost are non-negative), the sequence $\{\mathcal{L}(\Phi^{(t)})\}$ converges.
+By convexity, continuity, and compactness of $\mathcal{S}$, the iterates $\Phi^{(t)}$ converge to the unique minimizer $\Phi^{*}$.
+Now consider any initial distribution $p(\Phi^{(0)})$.
+Under the deterministic mapping $\mathcal{F}$, the sequence $p(\Phi^{(t)})$ is given by: $p(\Phi^{(t)})=\delta(\Phi^{(t)})$ and since $\Phi^{(t)}\rightarrow\Phi^{*}$, the sequence of delta distributions $\delta(\Phi^{(t)})$ converges weakly to $\delta(\Phi^{*})$.
+That is, $p(\Phi^{(t)})\xrightarrow{w}\delta(\Phi^{*})$.
 Appendix F Proof of Theorem 6